### 05 Mar 2026 v1

# Stochastic Particle Acceleration during Pressure-Anisotropy-Driven Magnetogenesis in the Pre-Structure Universe

Abstract
We investigate whether stochastic acceleration associated with pressure-anisotropy-driven magnetogenesis can generate a dynamically significant population of cosmic rays (CRs) prior to nonlinear structure formation. As magnetic fields amplify in the early Universe, the associated increase in gyrofrequency enhances pitch-angle scattering, potentially shortening the stochastic acceleration time. We derive an analytic criterion for efficient cosmological acceleration by comparing the acceleration timescale with the Hubble time, which defines a critical magnetic field and a corresponding CR turn-on redshift $z_{\rm on}$. For representative parameters, we find $z_{\rm on}\sim 1.7$. To quantify the resulting particle population, we solve a Fokker-Planck equation for the isotropic proton distribution in the redshift interval $z=10\rightarrow z_{\rm on}$. Throughout most of this epoch, adiabatic expansion dominates over stochastic energization and the distribution remains close to a cooling Maxwellian. However, as the system approaches the turn-on epoch, the stochastic acceleration time decreases, allowing a mild suprathermal tail to develop. Even under optimistic assumptions corresponding to the strong-scattering limit, the maximum attainable proton energy reaches at most $\mathcal{O}(10^{2})\,\mathrm{GeV}$. These results indicate that efficient CR production in the intergalactic medium is intrinsically tied to the onset of structure-formation shocks, while earlier microinstability-driven stochastic processes can provide at most a modest pre-acceleration.
I Introduction
Cosmic rays (CRs) are a ubiquitous nonthermal component of the Universe, observed across environments ranging from galaxies to galaxy clusters. In the Milky Way, the CR energy density is of order $\sim 1~{\rm eV\,cm^{-3}}$, comparable to that of the thermal interstellar medium, implying that CRs play a dynamically significant role in galactic systems [e.g., 21]. It is widely accepted that supernova remnants accelerate CRs up to $\sim 10^{15}$ eV via diffusive shock acceleration (DSA) [e.g., 2, 3, 5], and multiwavelength observations provide strong evidence for relativistic particle acceleration at such shocks [e.g., 16, 22, 1]. In some cases, CR precursors predicted by DSA have been observationally identified [24]. In the context of large-scale structure formation, DSA at collisionless shocks is likewise regarded as a primary mechanism responsible for accelerating particles to relativistic energies. Structure-formation shocks with velocities $u_{s}\sim 10^{2}$$10^{3}~{\rm km\,s^{-1}}$ naturally arise during hierarchical clustering, and numerical simulations support their role in generating CR populations in the intracluster and intergalactic media [e.g., 20, 28, 26, 14, 33, 34, 15, 29, 9, 10, 11].
However, an important open question concerns the origin of the first nonthermal particles. Cosmological simulations indicate that a large fraction of gas thermalization and cosmic-ray production occurs at structure-formation shocks emerging at relatively low redshifts ($z\lesssim{\rm a~few}$) [e.g., 28]. If CR production is tied exclusively to such shocks, then a pre-existing suprathermal seed population may not be required. Nevertheless, energetic events at earlier epochs may still generate the first CRs. For instance, early shock-driven scenarios associated with the first supernova explosions at $z\sim 20$ have been proposed as possible sources [e.g., 19, 23]. Another possibility arises during the subsequent pre-structure epoch, before large-scale structure formation enters the nonlinear regime and cosmological shocks become widespread ($z\lesssim\text{a few}$). During this period, the Universe undergoes a prolonged phase of weak magnetization, and magnetic fields are expected to grow gradually through various amplification processes [e.g., 27, 35]. If weak but cumulative acceleration processes operate during this intermediate redshift interval ($z\sim 10\rightarrow{\rm a~few}$), they could establish a low-level CR background that modifies subsequent injection and acceleration at structure-formation shocks. Determining whether such a pre-shock population exists is therefore relevant for understanding both the efficiency of DSA and the thermal history of the early intergalactic medium.
In parallel, recent studies of cosmological magnetogenesis have emphasized the role of pressure-anisotropy-driven plasma instabilities in amplifying magnetic fields in weakly magnetized environments [e.g., 30, 31, 32, 6, 12]. As the magnetic field grows, the associated proton gyrofrequency increases, enhancing pitch-angle scattering and potentially reducing the characteristic acceleration time of stochastic (second-order Fermi) processes operating in turbulence [e.g., 7, 4, 25]. This raises a natural question: can instability-assisted stochastic acceleration during cosmological magnetogenesis generate a dynamically significant CR population prior to the emergence of structure-formation shocks?
In this work, we investigate this possibility by combining (i) an analytic acceleration-time criterion based on the comparison between the stochastic acceleration time and the Hubble time, and (ii) a Fokker-Planck model for the isotropic proton distribution. This approach allows us to quantify both the redshift at which stochastic acceleration becomes cosmologically viable and the magnitude of any resulting nonthermal population. The central question we address is whether instability-enhanced scattering during cosmological magnetogenesis can reduce the stochastic acceleration time below the cosmological expansion timescale over any redshift interval prior to the widespread emergence of structure-formation shocks. In particular, we derive a critical condition for cosmologically viable second-order Fermi acceleration and follow the redshift evolution of the proton distribution under the combined effects of stochastic diffusion and adiabatic expansion. This framework enables a self-consistent assessment of (i) the epoch at which stochastic acceleration may become competitive with cosmic expansion, and (ii) the magnitude of any resulting suprathermal population.
The paper is organized as follows. In Section II, we derive the stochastic acceleration timescale in the instability-mediated magnetogenesis framework and obtain an analytic criterion for the turn-on of cosmological acceleration. In Section III, we solve a minimal Fokker-Planck equation to quantify the pre-structure stochastic heating of the proton distribution. In Section IV, we compare the stochastic acceleration timescale with DSA in the Bohm-diffusion limit. Section V summarizes our results and discusses their cosmological implications.
II Stochastic Acceleration During Pressure-Anisotropy-Driven Magnetogenesis
Magnetic-field amplification mediated by pressure-anisotropy-driven instabilities can substantially modify particle transport properties in weakly magnetized plasmas. Since magnetic fields play a central role in modulating particle scattering, the growth of cosmic magnetic fields prior to the onset of large-scale structure formation may naturally provide conditions for stochastic particle acceleration. In this section, we examine whether particles can be accelerated on cosmological timescales through second-order Fermi processes operating in turbulence generated during magnetogenesis.
II.1 Acceleration Timescale
We adopt a stochastic acceleration scenario in which particles gain energy through repeated scattering off moving magnetic irregularities. The characteristic acceleration timescale can be written as
$t_{\rm acc}(E,z)\approx\left(\frac{c}{v_{\rm tur}(z)}\right)^{2}\frac{\lambda(E,z)}{c},$ (1)
where $v_{\rm tur}(z)$ is the turbulent velocity at the outer scale, $\lambda(E,z)$ is the particle mean free path, and $c$ is the speed of light. For relativistic particles, the mean free path is determined by the effective scattering frequency,
$\lambda(E,z)\approx\frac{v(E)}{\nu_{\rm eff}(z)}.$ (2)
The effective scattering rate is defined as
$\nu_{\rm eff}(z)\equiv\nu_{ii}(z)+\nu_{\rm scatt}(z),$ (3)
where $\nu_{ii}$ is the Coulomb collision frequency and $\nu_{\rm scatt}$ represents the scattering induced by pressure-anisotropy-driven instabilities.
Following the instability-mediated magnetogenesis framework [e.g., 12], the scattering frequency in the regime where plasma instabilities dominate can be expressed as
$\nu_{\rm scatt}\sim\left(|\Delta|-2\beta^{-1}\right)^{3/2}\omega_{\rm cp},$ (4)
where $\Delta$ is the pressure anisotropy factor, $\beta$ is the plasma beta parameter, and $\omega_{\rm cp}=eB/m_{p}c$ is the proton gyrofrequency associated with the evolving magnetic field. Substituting Eq. (2) into Eq. (1), the acceleration timescale becomes
$t_{\rm acc}(z)\approx\left(\frac{c}{v_{\rm tur}(z)}\right)^{2}\frac{v(E)}{c\nu_{\rm eff}(z)}.$ (5)
This expression highlights that magnetic-field amplification, through its impact on $\omega_{\rm cp}$ and $\nu_{\rm scatt}$, can significantly shorten the acceleration time. Throughout this work, we adopt the scattering-dominated regime, consistent with the transition discussed in previous studies [6, 12]. In the very early stages of magnetic-field amplification, collisional effects can in principle influence the plasma dynamics. However, as the magnetic field grows and the proton gyrofrequency $\omega_{\rm cp}\propto B$ increases, the ratio $\nu_{\rm coll}/\omega_{\rm cp}\propto(1+z)^{3}B^{-1}$ rapidly decreases. For redshifts $z\lesssim 10$, where the baryon density has already declined substantially and magnetic amplification has progressed, the condition $\nu_{\rm coll}/\omega_{\rm cp}\ll 1$ is naturally satisfied. In this regime, pitch-angle scattering mediated by plasma instabilities dominates over Coulomb collisions ($\nu_{ii}\ll\nu_{\rm scatt}$). The effective scattering rate therefore satisfies $\nu_{\rm eff}\approx\nu_{\rm scatt}$, making the stochastic acceleration framework adopted here self-consistent.
To determine whether stochastic acceleration is viable before the formation of large-scale shocks, we compare the acceleration timescale with the Hubble time,
$t_{H}(z)=H^{-1}(z),$ (6)
where
$H(z)=H_{0}\sqrt{\Omega_{m}(1+z)^{3}+\Omega_{\Lambda}}.$ (7)
A necessary condition for efficient cosmological acceleration is therefore
$t_{\rm acc}(E,z)<t_{H}(z).$ (8)
Combining the above expressions yields the dimensionless ratio
$\frac{t_{\rm acc}}{t_{H}}\approx\left(\frac{c}{v_{\rm tur}}\right)^{2}\left(\frac{v(E)}{c}\right)\frac{H(z)}{\nu_{\rm eff}(z)}.$ (9)
When this ratio falls below unity, particles can be accelerated to relativistic energies within a cosmological expansion time. Conversely, if $t_{\rm acc}>t_{H}$, the expansion of the universe limits the achievable particle energies.
II.2 Implications for Pre-Shock CR Production
Eqs. (1)-(9) suggest that instability-driven scattering during magnetogenesis may create a channel for CR production prior to structure formation shocks. As the magnetic field grows, the associated increase in gyrofrequency enhances the scattering rate, thereby reducing the acceleration timescale. This feedback implies that even moderately amplified magnetic fields could enable stochastic acceleration in the early universe.
In this framework, the key quantity controlling the onset of acceleration is the ratio $H(z)/\nu_{\rm eff}(z)$. If instability-mediated scattering becomes sufficiently strong, stochastic acceleration may operate well before the emergence of collisionless shocks, potentially establishing a seed population of CRs that later participate in DSA. The results presented here therefore motivate a quantitative exploration of particle acceleration during cosmological magnetogenesis and its implications for the origin of the earliest nonthermal particle populations.
To evaluate the viability of stochastic acceleration during pressure-anisotropy-driven magnetogenesis, we adopt a minimal self-consistent closure of the background plasma parameters. Throughout this section we restrict our analysis to protons; extension to electrons will be discussed in future work.
Turbulent velocity.
We parametrize the turbulent velocity as a fixed fraction of the sound speed,
$v_{\rm tur}(z)=\epsilon\,c_{s}(z),$ (10)
where $\epsilon<1$ is a dimensionless parameter and
$c_{s}(z)=\sqrt{\frac{\gamma k_{B}T(z)}{\mu m_{p}}}$ (11)
is the adiabatic sound speed of the background plasma. We adopt $\mu\approx 0.59$ appropriate for a fully ionized primordial plasma. Observations and simulations of weakly magnetized astrophysical plasmas, such as the intracluster medium, suggest that gas motions are typically subsonic relative to the sound speed. For example, the Hitomi satellite measured a velocity dispersion of $\sim 160\,{\rm km\,s^{-1}}$ in the Perseus cluster core, indicating subsonic gas motions [13], while numerical studies likewise find turbulent velocities that are a fraction of the sound speed [8]. Motivated by these results, we explore a representative range $\epsilon=0.05$$0.5$.
Pressure anisotropy and plasma beta.
We assume that the pressure anisotropy remains near a quasi-stable value
$\Delta(z)\approx\Delta_{0}=\text{const}.$ (12)
Near the marginal stability condition, the plasma beta satisfies
$\beta(z)\sim\frac{2}{|\Delta_{0}|},$ (13)
which provides a closure relation between $\beta$ and the anisotropy amplitude. Taking into account the system for particle acceleration, the anisotropy factor of the system $\Delta=\Delta_{0}+\delta\Delta$ satisfies $|\Delta|>2\beta^{-1}$, where $\delta\Delta\ll\Delta_{0}$ represents a perturbation around the quasi-stable value $\Delta_{0}$, which is treated as a free parameter that drives the instability. The supercritical excess above the marginal threshold $|\Delta_{0}|\sim 2\beta^{-1}$ is
$\Delta_{\rm ex}(\delta,\beta)\equiv|\Delta_{0}+\delta\Delta|-|\Delta_{0}|=|\Delta_{0}+\delta\Delta|-2\beta^{-1},$ (14)
which reduces to $\Delta_{\rm ex}=\delta\Delta$ for $\Delta_{0}>0$ and $\delta\Delta\geq 0$. In high-$\beta$ astrophysical plasmas such as the intracluster medium (ICM), where $\beta\sim\mathcal{O}(10^{2})$ [e.g., 30, 31, 32, 27, 17, 18], this corresponds to a typical anisotropy level $\Delta_{0}\sim\mathcal{O}(10^{-2})$.
Magnetic-field evolution.
The magnetic field evolves according to the instability-driven growth equation derived in the scattering-dominated regime [12],
$\frac{dB}{dt}$ $\approx\omega_{\rm cp,0}\frac{\left(|\Delta|-2\beta^{-1}\right)^{3/2}}{1+|\Delta|}\left(\frac{B}{B_{0}}\right)^{2}$
$+\omega_{\rm cp,0}\frac{|\delta\Delta|^{3/2}}{1+|\Delta|}e^{-\Gamma_{d}t-\frac{1}{2}\Gamma_{d}/\Gamma_{c}^{(\rm inst)}}\left(\frac{B}{B_{0}}\right)^{2},$ (15)
where the second term represents the contribution from perturbations of the pressure anisotropy. Here $\Gamma_{d}$ denotes the effective damping rate of the anisotropy due to pitch-angle scattering by plasma instabilities, whereas $\Gamma_{c}^{(\rm inst)}$ represents the characteristic rate at which large-scale plasma motions generate pressure anisotropy. $B_{0}$ is the seed magnetic field and $\omega_{\rm cp,0}$ corresponds to the proton gyrofrequency evaluated at $B_{0}$.
In high-$\beta$ weakly magnetized plasmas, pressure anisotropy is expected to be regulated near the instability threshold by rapid pitch-angle scattering, so that the strong-damping limit ($\Gamma_{d}\gg\Gamma_{c}^{(\rm inst)}$) provides an appropriate closure. For small perturbations of the anisotropy ($|\delta\Delta/\Delta|\ll 1$), the perturbation-driven contribution becomes negligible. While the detailed magnetic-field growth timescale can depend on the relative magnitude of the damping and perturbation terms [12], these effects primarily modify the growth rate without significantly altering the overall magnetic-field evolution. The magnetic-field evolution therefore reduces to
$\frac{dB}{dt}\approx\omega_{\rm cp,0}\frac{\left(|\Delta|-2\beta^{-1}\right)^{3/2}}{1+|\Delta|}\left(\frac{B}{B_{0}}\right)^{2}.$ (16)
This nonlinear growth implies that once the instability condition $|\Delta|>2\beta^{-1}$ is satisfied, the magnetic field amplifies superlinearly with $B$, leading to rapid amplification during the magnetogenesis phase.
II.3 Analytic constraint on the critical magnetic field
A necessary condition for efficient cosmological acceleration is that the acceleration time be shorter than the Hubble time, $t_{\rm acc}<t_{H}=H^{-1}(z)$. Combining Eqs.  (4), (9), and (14) yields an analytic lower bound on the magnetic field:
$B>B_{\rm crit}(z)\equiv\frac{m_{p}c}{e}\left(\frac{c}{v_{\rm tur}(z)}\right)^{2}\left(\frac{v(E)}{c}\right)\frac{H(z)}{\Delta_{\rm ex}^{3/2}}.$ (17)
Eq. (17) makes explicit the strong parametric dependence $B_{\rm crit}\propto H(z)\,\epsilon^{-2}c_{s}^{-2}(z)\Delta_{\rm ex}^{-3/2}$: unless the turbulent motions are sufficiently strong and/or the anisotropy is strongly supercritical, the magnetic field required for $t_{\rm acc}<t_{H}$ becomes unrealistically large, implying that pressure-anisotropy-driven instabilities alone are unlikely to produce a significant pre-shock CR population under generic cosmological conditions.
Using Eq. (17), we can derive an explicit redshift dependence of the critical field required for $t_{\rm acc}<t_{H}$. At $z\gtrsim 1$ (matter-dominated era) we approximate
$H(z)\simeq H_{0}\sqrt{\Omega_{m}}\,(1+z)^{3/2}.$ (18)
Adopting a parametrized thermal history,
$T(z)=T_{0}\left(\frac{1+z}{1+z_{0}}\right)^{\alpha},$ (19)
the sound speed satisfies $c_{s}^{2}(z)\propto T(z)\propto(1+z)^{\alpha}$. Eq. (17) then yields
$B_{\rm crit}(z)\propto\frac{H(z)}{c_{s}^{2}(z)}\propto(1+z)^{\frac{3}{2}-\alpha},$ (20)
or equivalently
$B_{\rm crit}(z)=B_{\rm crit}(z_{0})\,\left(\frac{1+z}{1+z_{0}}\right)^{\frac{3}{2}-\alpha},$ (21)
where $B_{\rm crit}(z_{0})$ collects the normalization factors, including $\epsilon^{-2}$ and $\Delta_{\rm ex}^{-3/2}$. For the illustrative case $\alpha=1$ used in our toy thermal model, $B_{\rm crit}(z)\propto(1+z)^{1/2}$, implying a slowly increasing critical field toward higher redshift.
Figure 1: Redshift evolution of the magnetic field $B(z)$ (black) compared with the analytic critical field $B_{\rm crit}(z)$ (red). Their intersection defines the CR turn-on redshift $z_{\rm on}$. The curves are shown for fiducial parameters $B_{0}=10^{-18}$ G, $\delta\Delta/\Delta_{0}=0.1$, $\Delta_{0}\sim 2\beta^{-1}=10^{-2}$, and $\epsilon=0.2$.
Figure 2: Dependence of the CR turn-on redshift $z_{\rm on}$ on (a) the turbulence amplitude $\epsilon$ and (b) the pressure-anisotropy perturbation $\delta\Delta$. All other parameters are fixed to the fiducial values used in Fig. 1. Over the explored parameter range, $z_{\rm on}$ varies only weakly, remaining close to $z_{\rm on}\sim 1.7$.
II.4 Definition of the CR turn-on redshift
A useful way to quantify when pressure-anisotropy-driven scattering can begin to enable cosmological stochastic acceleration is to define a “CR turn-on redshift” $z_{\rm on}$. We define $z_{\rm on}$ as the epoch at which the stochastic acceleration time becomes comparable to the Hubble time:
$\left.\frac{t_{\rm acc}(z)}{t_{H}(z)}\right|_{z=z_{\rm on}}=1,\qquad t_{H}(z)=H^{-1}(z).$ (22)
For $z>z_{\rm on}$ the expansion time is too short ($t_{\rm acc}>t_{H}$) and stochastic acceleration is inefficient, while for $z<z_{\rm on}$ acceleration becomes viable ($t_{\rm acc}<t_{H}$).
Using the analytic constraint derived above, the turn-on condition can be written equivalently as a crossing between the evolving magnetic field and the critical field:
$B(z_{\rm on})=B_{\rm crit}(z_{\rm on}),$ (23)
where $B_{\rm crit}(z)$ is defined by Eq. (17). In the matter-dominated regime $H(z)\propto(1+z)^{3/2}$ and for a parametrized thermal history $T(z)\propto(1+z)^{\alpha}$ (so that $c_{s}^{2}\propto(1+z)^{\alpha}$), one finds $B_{\rm crit}(z)\propto(1+z)^{3/2-\alpha}$.
For illustrative purposes, if the magnetic-field evolution can be approximated by $B(z)=B_{0}(1+z)^{p}$, the turn-on redshift follows from Eq. (23) as
$1+z_{\rm on}=\left(\frac{C}{B_{0}}\right)^{\!\frac{1}{\,p-\left(\frac{3}{2}-\alpha\right)}},$ (24)
where $C$ collects the normalization factors in $B_{\rm crit}$, including the strong dependences $C\propto\epsilon^{-2}\Delta_{\rm ex}^{-3/2}$. A finite $z_{\rm on}$ therefore exists only if the magnetic field grows sufficiently rapidly with redshift, $p>(3/2-\alpha)$; otherwise the system remains in the acceleration-inefficient regime at all epochs.
To illustrate this behavior explicitly, we compute the magnetic-field evolution $B(z)$ using the unperturbed scattering-dominated solution from our previous work and compare it with the analytic critical field $B_{\rm crit}(z)$. Fig. 1 shows the resulting redshift evolution for a representative set of parameters. At high redshift the magnetic field remains far below the critical value, implying $t_{\rm acc}\gg t_{H}$ and therefore inefficient stochastic acceleration. As cosmic time progresses, nonlinear magnetic amplification sets in and $B(z)$ increases rapidly. The CR turn-on redshift $z_{\rm on}$ is defined by the crossing $B(z)=B_{\rm crit}(z)$. In the fiducial case shown here, this occurs at $z_{\rm on}\sim 1.7$, corresponding to a magnetic field strength of order $10^{-9}$ G. Importantly, this epoch coincides with the onset of nonlinear structure formation, suggesting that efficient CR production is naturally associated with structure-formation shocks rather than with earlier microinstability-driven stochastic processes.
Fig. 2 illustrates the dependence of the turn-on redshift on the turbulence amplitude $\epsilon$ and on the pressure anisotropy, parameterized by $\delta\Delta/\Delta_{0}$ and $\Delta_{0}\sim 2\beta^{-1}$. Despite the explicit appearance of both quantities in the analytic expression for $B_{\rm crit}(z)$, the resulting variation in $z_{\rm on}$ is remarkably small over the considered parameter range. In particular, even when $\epsilon$ and $\delta\Delta$ are varied by factors of a few, the turn-on redshift remains confined to a narrow interval around $z_{\rm on}\sim 1.7$. This behavior reflects the fact that $B(z)$ exhibits a rapid nonlinear growth phase, so that modest vertical shifts in $B_{\rm crit}(z)$ translate into only small horizontal shifts in the crossing redshift. The onset of cosmologically viable stochastic acceleration is therefore robust against moderate changes in the underlying microphysical parameters. Importantly, the inferred $z_{\rm on}$ remains close to the epoch of nonlinear structure formation, reinforcing the conclusion that efficient CR production is naturally associated with structure-formation shocks rather than with earlier microinstability-driven processes.
Figure 3: Redshift evolution of the maximum attainable proton energy $E_{\max}(z)$ computed from Eq. (28) for turbulence amplitudes $\epsilon=0.05$ (black), $0.1$ (red), and $0.2$ (blue). Near the turn-on epoch $z_{\rm on}\sim 1.7$, $E_{\max}$ reaches $\sim 10^{2}\,\mathrm{GeV}$.
II.5 Maximum attainable energy during cosmological magnetogenesis
While the condition $t_{\rm acc}(z)<t_{H}(z)$ determines when stochastic acceleration becomes cosmologically viable, it does not specify the maximum particle energy that can be reached. To quantify the physical significance of the turn-on epoch, we estimate the maximum proton energy attainable before the expansion of the Universe limits further acceleration.
Acceleration time: optimistic upper bound.
To estimate the largest particle energies that could in principle be reached, we adopt an optimistic upper bound for the particle mean free path by assuming $\lambda\sim r_{L}$, where $r_{L}$ is the Larmor radius. This corresponds to the strong-scattering limit $\nu_{\rm scatt}\sim\omega_{\rm cp}$, which can arise when instability-driven pitch-angle scattering dominates. In this limit, the spatial diffusion coefficient becomes
$D(E)\sim\frac{1}{3}r_{L}(E)c,\qquad r_{L}(E)=\frac{pc}{eB}\approx\frac{E}{eB},$ (25)
for relativistic protons with $E\approx pc$. In this limit, the second-order Fermi acceleration time can be expressed as
$t_{\rm acc,2}(E,z)\approx\left(\frac{c}{v_{\rm tur}(z)}\right)^{2}\frac{r_{L}(E)}{c}=\left(\frac{c}{v_{\rm tur}(z)}\right)^{2}\frac{E}{eB(z)c}.$ (26)
This expression captures the essential energy dependence: higher-energy particles require longer acceleration times.
Definition of the maximum energy.
The maximum attainable energy at a given redshift is obtained by requiring that the acceleration time not exceed the Hubble time,
$t_{\rm acc,2}(E_{\max},z)=t_{H}(z)=H^{-1}(z).$ (27)
Substituting the parametrization $v_{\rm tur}(z)=\epsilon c_{s}(z)$ into Eq. (26), the maximum attainable energy becomes
$E_{\max}(z)\approx eB(z)c\epsilon^{2}\left(\frac{c_{s}(z)}{c}\right)^{2}H^{-1}(z).$ (28)
In the matter-dominated regime, $H(z)\propto(1+z)^{3/2}$. Adopting the parametrized thermal history $T(z)\propto(1+z)^{\alpha}$, so that $c_{s}^{2}(z)\propto(1+z)^{\alpha}$, we find
$E_{\max}(z)\propto B(z)\,(1+z)^{\alpha-3/2}.$ (29)
Thus, even if the magnetic field grows rapidly, cosmological expansion suppresses the maximum achievable energy at high redshift.
Physical implications.
Fig. 3 shows the redshift evolution of the maximum attainable proton energy for representative values of the turbulence parameter $\epsilon$. Evaluated near the turn-on epoch $z_{\rm on}\sim 1.7$, where $B(z_{\rm on})\sim 10^{-9}\,\mathrm{G}$ in our fiducial model, Eq. (28) implies that proton energies can, under optimistic optimistic assumptions corresponding to the strong-scattering limit, reach values up to $\sim 10^{2}\,\mathrm{GeV}$. The strong dependence $E_{\max}\propto\epsilon^{2}$ implies that larger turbulence amplitudes lead to higher maximum energies, while the overall redshift dependence remains unchanged. At higher redshift ($z\gtrsim z_{\rm on}$), the magnetic field remains weak and the Hubble time short, so that $E_{\max}$ is strongly suppressed and remains far below the relativistic regime. Thus, even under optimistic assumptions, cosmological stochastic acceleration cannot generate a substantial high-energy particle population prior to the onset of nonlinear structure formation. We emphasize that Eq. (28) provides an upper bound. Additional physical effects—including Coulomb and ionization losses, incomplete turbulence development, or deviations from Bohm-like diffusion—would further reduce the achievable energy. Consequently, while stochastic acceleration may in principle produce $\mathcal{O}(10$-$10^{2})\,\mathrm{GeV}$ protons near $z_{\rm on}\sim 1.7$, the emergence of structure-formation shocks remains a far more efficient and robust mechanism for accelerating CRs to higher energies.
III Pre-Structure Stochastic Acceleration in the Early Universe
To quantify the magnitude of any seed nonthermal population produced by the instability-assisted stochastic process discussed above, we introduce a transport model for the isotropic proton distribution function. The goal is not to construct a fully self-consistent cosmological CR population synthesis, but rather to translate the acceleration-time constraints into an explicit estimate of the resulting momentum-space spectrum and energy content.
III.1 Isotropic Fokker-Planck equation in an expanding background
Let $f(p,z)$ be the isotropic phase-space distribution of protons, normalized such that the differential number density is $dn=4\pi p^{2}f(p,z)\,dp$. Neglecting spatial gradients and assuming isotropic pitch-angle scattering, the momentum-space Fokker-Planck equation reads
$\frac{\partial f}{\partial t}$ $=\frac{1}{p^{2}}\frac{\partial}{\partial p}\left[p^{2}D_{pp}(p,z)\,\frac{\partial f}{\partial p}\right]$
$-\frac{1}{p^{2}}\frac{\partial}{\partial p}\left[p^{2}\,\dot{p}_{\rm loss}(p,z)\,f\right]+Q(p,z)-\frac{f}{t_{\rm esc}(p,z)},\,$ (30)
where $D_{pp}$ is the momentum diffusion coefficient associated with second-order Fermi acceleration, $\dot{p}_{\rm loss}$ represents systematic momentum losses, $Q$ is a source term (e.g., injection from the thermal pool), and $t_{\rm esc}$ is an effective escape time from the acceleration region.
Cosmological expansion is incorporated by transforming from $t$ to redshift $z$ using
$\frac{dt}{dz}=-\frac{1}{(1+z)H(z)}\,$, (31)
with $H(z)=H_{0}\sqrt{\Omega_{\rm m}(1+z)^{3}+\Omega_{\Lambda}}$. Eq. (30) then becomes
$-(1+z)H(z)\,\frac{\partial f}{\partial z}$ $=\frac{1}{p^{2}}\frac{\partial}{\partial p}\left[p^{2}D_{pp}(p,z)\,\frac{\partial f}{\partial p}\right]$
$-\frac{1}{p^{2}}\frac{\partial}{\partial p}\left[p^{2}\,\dot{p}_{\rm loss}(p,z)\,f\right]+Q(p,z)\,.$ (32)
In a second-order Fermi process associated with the pressure-anisotropy-driven instabilities, which generate small-scale magnetic fluctuations that scatter particles, the characteristic acceleration time can be expressed in terms of $D_{pp}$ as
$D_{pp}(p,z)=\frac{p^{2}}{t_{\rm acc}(z)}~(p\gtrsim p_{\rm inj}),$ (33)
corresponding to a momentum diffusion coefficient proportional to $p^{2}$, with a redshift-dependent normalization determined by the microphysical scattering rate. Here, $p_{\rm inj}$ denotes the injection momentum for stochastic acceleration. This form represents the minimal choice consistent with the timescale analysis presented in Section II.
In the present model, we assume that the escape timescale is much longer than the characteristic acceleration and loss timescales, $t_{\rm esc}\gg t_{\rm acc},~t_{\rm loss}$. Here, the characteristic loss timescale is defined as $t_{\rm loss}\equiv p/\dot{p}_{\rm loss}$. Cosmological expansion inevitably induces adiabatic momentum losses,
$\dot{p}_{\rm ad}(p,z)=-H(z)\,p\,.$ (34)
In principle, additional Coulomb losses can be included at sub-relativistic energies, $\dot{p}_{\rm loss}=\dot{p}_{\rm ad}+\dot{p}_{\rm C}$, to assess whether any incipient suprathermal tail is efficiently thermalized. However, since our main interest lies in the relativistic tail near and above the turn-on epoch, the adiabatic term already captures the unavoidable cosmological sink. Therefore, we adopt $\dot{p}_{\rm loss}\approx\dot{p}_{\rm ad}$. With Eqs. (33) and (34), Eq. (32) becomes
$-(1+z)H(z)\,\frac{\partial f}{\partial z}$ $\approx\frac{1}{p^{2}}\frac{\partial}{\partial p}\left[p^{2}\frac{p^{2}}{t_{\rm acc}(z)}\,\frac{\partial f}{\partial p}\right]$
$-\frac{1}{p^{2}}\frac{\partial}{\partial p}\left[p^{2}\left(-H(z)p\right)f\right]+Q(p,z)\,.$ (35)
Figure 4: Proton spectra $p^{4}f(p,z)$ from the Fokker-Planck model for $z=10,5,2$ and $1.7$. Panel (a): $\epsilon=0.05$; panel (b): $\epsilon=0.2$. The initial distribution at $z=10$ is Maxwellian with $k_{B}T_{0}=0.86~{\rm eV}$.
Figure 5: Same as Fig. 4, but including a weak pre-existing suprathermal power-law tail ($s=4.2$) in the initial distribution.
III.2 Numerical implementation and early-universe heating
To assess the magnitude of stochastic heating prior to the magnetic-field turn-on epoch, we solve Eq. (35) numerically in the redshift interval $z=10\rightarrow z_{\rm on}$, where the magnetic field remains well below its rapid-amplification regime. In this early phase, the goal is to determine whether the instability-assisted second-order Fermi process can produce a significant seed nonthermal component over a Hubble time.
The initial condition at $z=10$ is taken to be a non-relativistic Maxwellian distribution,
$f_{\rm MW}(p,z=10)=\frac{n}{(2\pi m_{p}k_{B}T_{0})^{3/2}}\exp\!\left(-\frac{p^{2}}{2m_{p}k_{B}T_{0}}\right),$ (36)
where $n$ denotes the proton number density and $k_{B}T_{0}=0.86~{\rm eV}$. No explicit source term is included ($Q=0$), and escape is neglected, so that the evolution is governed solely by stochastic momentum diffusion and adiabatic losses. The magnetic field $B(z)$ is mapped to redshift using the cosmological relation $dt/dz=-[(1+z)H(z)]^{-1}$. In the early regime ($z\gtrsim z_{\rm on}$), $B(z)$ evolves smoothly and remains far below the explosive growth associated with $z<z_{\rm on}$, ensuring that $t_{\rm acc}(z)$ varies gradually.
Eq. (35) is solved on a logarithmic momentum grid using a fully implicit scheme in redshift. The equation is written in conservative flux form in momentum space,
$\mathcal{R}(f)=\frac{1}{p^{2}}\frac{\partial}{\partial p}\left[p^{2}D_{pp}\frac{\partial f}{\partial p}-p^{2}\dot{p}_{\rm ad}f\right],$ (37)
and discretized such that particle number conservation is maintained up to numerical accuracy. The redshift step $\Delta z$ is mapped to an effective time step $\Delta t=-\Delta z/[(1+z)H(z)]$, ensuring consistency with the cosmological evolution.
Fig. 4 shows the resulting spectra $p^{4}f(p,z)$ at $z=10,5,$ and $2$ for two representative turbulence levels, $\epsilon=0.05$ and $\epsilon=0.2$. In both cases, the thermal peak shifts toward lower momentum as redshift decreases. This behavior reflects adiabatic cooling, for which $p\propto a^{-1}\propto(1+z)$. The corresponding reduction in peak amplitude is likewise consistent with the decreasing thermal energy density of the expanding background. The stochastic term introduces a modest broadening of the high-momentum tail relative to pure adiabatic evolution. For $\epsilon=0.05$, the acceleration time remains long compared to the Hubble time throughout $z=10\rightarrow 2$, and the distribution remains close to Maxwellian. For $\epsilon=0.2$, the diffusion coefficient increases by a factor of four, leading to a slightly enhanced suprathermal component; however, adiabatic cooling still dominates the overall evolution over this redshift interval.
As the system approaches the magnetic-field turn-on epoch ($z\sim z_{\rm on}\sim 1.7$), the situation begins to change. The amplification of the magnetic field reduces the stochastic acceleration time $t_{\rm acc}$, allowing the diffusion term to partially compete with the adiabatic loss term. As a result, a mild suprathermal tail begins to emerge in the particle spectrum. Although the nonthermal component remains energetically subdominant, this behavior indicates that stochastic acceleration becomes progressively more effective near the onset of magnetic-field amplification.
We do not extend the mini Fokker–Planck integration below $z_{\rm on}$, since in this regime large-scale structure formation shocks are expected to dominate particle acceleration. The present model is therefore designed to quantify the pre-shock stochastic heating phase and the magnitude of any seed suprathermal population prior to the onset of structure formation. In the next section, we compare the characteristic acceleration timescale of the instability-assisted stochastic process with that of DSA in the Bohm-diffusion limit, in order to determine which mechanism governs CR production once shocks emerge.
III.3 Sensitivity to a pre-existing suprathermal population
The calculations above assume that the proton distribution at $z=10$ is purely Maxwellian. However, weak nonthermal particle populations may already exist in the early universe as a result of earlier energetic events, such as Population III supernova explosions or structure-formation shocks [e.g., 23].
To assess the sensitivity of the stochastic heating process to such initial conditions, we consider an alternative initial distribution consisting of a Maxwellian core with a small suprathermal power-law tail,
$f(p,z=10)=f_{\rm MW}(p,z=10)+A\,p^{-s}\exp(-p/p_{\rm cut}).$ (38)
We then evolve this distribution using Eq. (35). The slope of the pre-existing tail is taken to be $s=4.2$, motivated by typical spectra expected from DSA. The normalization of the tail $A$ is chosen such that the nonthermal component matches a small fraction of the thermal Maxwellian at a reference momentum $p_{0}$, while an exponential cutoff $p_{\rm cut}$ limits the high-momentum extent of the seed population.
The resulting spectra are shown in Fig. 5. We find that the overall evolution from $z=10$ to $z=2$ is largely insensitive to the presence of this weak pre-existing suprathermal population. In both turbulence cases ($\epsilon=0.05$ and $\epsilon=0.2$), the proton distribution shifts primarily toward lower momenta as the universe expands. This behavior arises from the adiabatic momentum loss term $\dot{p}_{\rm ad}=-H(z)p$, which implies $p\propto a^{-1}\propto(1+z)$. Although stochastic momentum diffusion slightly broadens the high-momentum tail, particularly for larger turbulence amplitudes, adiabatic cooling remains the dominant influence on the spectral evolution over the interval $z=10\rightarrow 2$.
As the system approaches the magnetic-field turn-on epoch ($z\sim z_{\rm on}\sim 1.7$), the reduction of the stochastic acceleration timescale allows the diffusion term to partially compete with adiabatic losses. Consequently, a mild suprathermal tail begins to emerge in the particle spectrum. The presence of a pre-existing power-law tail does not significantly alter this behavior, indicating that the early evolution is largely controlled by cosmological cooling rather than by stochastic re-acceleration of the seed population.
IV Comparison with shock acceleration in the Bohm-diffusion limit
Even if stochastic acceleration becomes formally viable after the turn-on epoch $z_{\rm on}$, it is important to assess whether it can compete with acceleration in structure-formation shocks. To this end, we compare the characteristic acceleration timescale of our stochastic acceleration mechanism with the standard DSA timescale assuming Bohm-like diffusion.
DSA with Bohm-like diffusion.
For a non-relativistic shock of speed $u_{s}$ (in the upstream frame), the DSA acceleration time can be written as [e.g., 5]
$t_{\rm acc,sh}(E)\approx\frac{3}{u_{1}-u_{2}}\left(\frac{D_{1}(E)}{u_{1}}+\frac{D_{2}(E)}{u_{2}}\right),$ (39)
where $u_{1,2}$ and $D_{1,2}$ are the upstream/downstream flow speeds and spatial diffusion coefficients, respectively. For a strong shock with compression ratio $r=u_{1}/u_{2}\simeq 4$ and comparable diffusion coefficients on both sides ($D_{1}\sim D_{2}\sim D$), Eq. (39) reduces to the commonly used estimate
$t_{\rm acc,sh}(E)\sim\xi\,\frac{D(E)}{u_{s}^{2}},\qquad\xi\sim\mathcal{O}(10),$ (40)
where $\xi$ absorbs order-unity factors depending on the shock structure. In the Bohm-like limit, the diffusion coefficient is
$D_{\rm B}(E)\equiv\frac{1}{3}r_{L}(E)c,\qquad r_{L}(E)=\frac{pc}{eB}\approx\frac{E}{eB},$ (41)
for relativistic particles with $E\approx pc$. Combining Eqs. (40)–(41) yields
$t_{\rm acc,sh}(E)\sim\frac{\xi}{3}\,\frac{r_{L}(E)c}{u_{s}^{2}}\sim\frac{\xi}{3}\,\frac{E}{eB}\,\frac{c}{u_{s}^{2}}.$ (42)
Timescale ratio and dominance of shocks.
To compare the characteristic acceleration timescales of DSA and stochastic acceleration, we form the ratio
$\frac{t_{\rm acc,2}}{t_{\rm acc,sh}}\sim\frac{3}{\xi}\,\left(\frac{c}{v_{\rm tur}}\right)^{2}\left(\frac{u_{s}}{c}\right)^{2}\left(\frac{c}{r_{L}\,\nu_{\rm eff}}\right),$ (43)
where $\xi\sim\mathcal{O}(10)$ parameterizes order-unity factors in the DSA timescale. The final factor, $c/(r_{L}\nu_{\rm eff})$, measures the efficiency of pitch-angle scattering over a gyro-orbit; in the Bohm-like limit, $\nu_{\rm eff}\sim c/r_{L}$, this factor is of order unity.
For structure-formation environments, the relevant characteristic velocities are $v_{\rm tur}\sim\epsilon c_{s}$ with $c_{s}\sim 10^{2}-10^{3}~{\rm km\,s^{-1}}$ and $\epsilon\sim 0.1$, implying $v_{\rm tur}/c\sim 10^{-4}-10^{-3}$. Large-scale structure shocks typically have $u_{s}\sim 10^{3}~{\rm km\,s^{-1}}$, corresponding to $u_{s}/c\sim 3\times 10^{-3}$. Adopting Bohm-like scattering ($c/(r_{L}\nu_{\rm eff})\sim 1$), these scalings imply
$\frac{t_{\rm acc,2}}{t_{\rm acc,sh}}\sim 10-10^{3},$ (44)
up to order-unity factors. Thus, even under optimistic assumptions for the turbulence-driven stochastic mechanism, second-order acceleration is typically one to three orders of magnitude slower than shock acceleration in the same environment.
Consequently, once nonlinear structure formation generates shocks with $u_{s}\sim 10^{3}~{\rm km\,s^{-1}}$ and simultaneously amplifies the magnetic field, DSA operates on substantially shorter timescales. Even if the instability-assisted stochastic channel formally turns on at $z\lesssim z_{\rm on}$, shock acceleration is expected to dominate the production of CRs in that epoch. This supports our central conclusion: a significant pre-shock CR population is unlikely, and efficient CR production is naturally tied to the onset of structure-formation shocks.
V Summary and Discussion
We investigated whether stochastic (second-order Fermi) acceleration during pressure-anisotropy-driven magnetogenesis can generate a dynamically significant CR population prior to nonlinear structure formation. By combining analytic acceleration-time constraints with a Fokker-Planck model, we quantified both the turn-on condition for efficient stochastic acceleration and the resulting proton momentum distribution in the early Universe. The analytic criterion $t_{\rm acc}<t_{H}$ yields a critical magnetic field $B_{\rm crit}(z)$ and defines a CR turn-on redshift $z_{\rm on}\sim 1.7$ for representative parameters. At higher redshift, the magnetic field remains below this threshold, so that stochastic acceleration proceeds too slowly to compete with cosmological expansion. Even near the turn-on epoch, the maximum attainable proton energy, evaluated under optimistic assumptions corresponding to the strong-scattering limit, reaches at most $\mathcal{O}(10^{2})\,\mathrm{GeV}$. The numerical Fokker-Planck calculation shows that, throughout $z=10\rightarrow z_{\rm on}$, the proton distribution remains close to a cooling Maxwellian. Adiabatic expansion dominates the evolution, and the resulting nonthermal energy density remains far below the thermal background energy density.
These results carry several cosmological implications. First, instability-mediated stochastic acceleration is unlikely to modify the thermal history of the pre-structure intergalactic medium. Second, the resulting suprathermal component is too weak to produce observable distortions of the cosmic microwave background or to contribute significantly to diffuse gamma-ray or neutrino backgrounds. Finally, efficient CR production appears to be intrinsically tied to the onset of nonlinear structure formation, when large-scale shocks and amplified magnetic fields operate on substantially shorter acceleration timescales. While the low-level suprathermal population found here may in principle serve as a seed for subsequent DSA, its energy content is modest. A detailed assessment of the injection efficiency and the reprocessing of this seed population by structure-formation shocks will require a dedicated kinetic calculation, which we defer to future work.
In summary, pressure-anisotropy-driven magnetogenesis can modulate particle scattering in the early Universe, but it does not by itself generate a dynamically important pre-shock CR population. The emergence of efficient CR acceleration is therefore naturally associated with the epoch of nonlinear structure formation.