### 01 Apr 2026 v2

# Stochastic Particle Acceleration during Pressure-Anisotropy-Driven Magnetogenesis in the Pre-Structure Universe

Abstract
We investigate whether stochastic acceleration associated with pressure-anisotropy-driven magnetogenesis can generate a dynamically significant population of cosmic rays (CRs) prior to nonlinear structure formation. As magnetic fields amplify in the early Universe, the associated increase in gyrofrequency enhances pitch-angle scattering, potentially shortening the stochastic acceleration time. We derive an analytic criterion for efficient cosmological acceleration by comparing the acceleration timescale with the Hubble time, which defines a critical magnetic field and a corresponding CR turn-on redshift $z_{\rm on}$. For representative parameters, we find $z_{\rm on}\sim 1.7$. To quantify the resulting particle population, we solve a Fokker-Planck equation for the isotropic ion (proton) distribution in the redshift interval $z=10\rightarrow z_{\rm on}$, including Coulomb energy losses in a fully ionized intergalactic medium. Throughout most of this epoch, adiabatic expansion dominates over stochastic energization, and Coulomb cooling efficiently thermalizes low-energy particles, introducing an effective injection threshold at energies of order ${\mathcal{O}}(10)$ keV. As a result, the distribution remains close to a cooling Maxwellian, and the formation of a suprathermal tail is strongly suppressed even in the presence of a pre-existing nonthermal component. Even under optimistic assumptions corresponding to the strong-scattering limit, the maximum attainable ion energy reaches at most $\mathcal{O}(10^{2})$ GeV. These results indicate that efficient CR production in the intergalactic medium is intrinsically tied to the onset of structure-formation shocks, while earlier microinstability-driven stochastic processes can provide at most a modest pre-acceleration.
1 Introduction
Cosmic rays (CRs) are ubiquitous nonthermal components of the Universe, observed across environments ranging from galaxies to galaxy clusters. In the Milky Way, the CR energy density is of order $\sim 1~{\rm eV\,cm^{-3}}$, comparable to that of the thermal interstellar medium, implying that CRs play a dynamically significant role in galactic systems [e.g., 22]. It is widely accepted that supernova remnants accelerate CRs up to $\sim 10^{15}$ eV via diffusive shock acceleration (DSA) [e.g., 2, 3, 5], and multiwavelength observations provide strong evidence for relativistic particle acceleration at such shocks [e.g., 17, 23, 1]. In some cases, CR precursors predicted by DSA have been observationally identified [25]. In the context of large-scale structure formation, DSA at collisionless shocks is likewise regarded as a primary mechanism responsible for accelerating particles to relativistic energies. Structure-formation shocks with velocities $u_{s}\sim 10^{2}$$10^{3}~{\rm km\,s^{-1}}$ naturally arise during hierarchical clustering, and numerical simulations support their role in generating CR populations in the intracluster and intergalactic media [e.g., 21, 29, 27, 15, 36, 37, 16, 31, 10, 11, 12].
However, an important open question concerns the origin of the first nonthermal particles. Cosmological simulations indicate that a large fraction of gas thermalization and cosmic-ray production occurs at structure-formation shocks emerging at relatively low redshifts ($z\lesssim{\rm a~few}$) [e.g., 29]. If CR production is tied exclusively to such shocks, then a pre-existing suprathermal seed population may not be required. Nevertheless, energetic events at earlier epochs may still generate the first CRs. For instance, early shock-driven scenarios associated with the first supernova explosions at $z\sim 20$ have been proposed as possible sources [e.g., 20, 24]. Another possibility arises during the subsequent pre-structure epoch, before large-scale structure formation enters the nonlinear regime and cosmological shocks become widespread ($z\lesssim\text{a few}$). During this period, the Universe undergoes a prolonged phase of weak magnetization, and magnetic fields are expected to grow gradually through various amplification processes [e.g., 28, 38]. If weak but cumulative acceleration processes operate during this intermediate redshift interval ($z\sim 10\rightarrow{\rm a~few}$), they could establish a low-level CR background that modifies subsequent injection and acceleration at structure-formation shocks. Determining whether such a pre-shock population exists is therefore relevant for understanding both the efficiency of DSA and the thermal history of the early intergalactic medium.
In parallel, recent studies of cosmological magnetogenesis have emphasized the role of pressure-anisotropy-driven plasma instabilities in amplifying magnetic fields in weakly magnetized environments [e.g., 32, 33, 34, 6, 13]. As the magnetic field grows, the associated ion (proton) gyrofrequency increases, enhancing pitch-angle scattering and potentially reducing the characteristic acceleration time of stochastic (second-order Fermi) processes operating in turbulence [e.g., 7, 4, 26]. This raises a natural question: can instability-assisted stochastic acceleration during cosmological magnetogenesis generate a dynamically significant CR population prior to the emergence of structure-formation shocks?
In this work, we investigate this possibility by combining (i) an analytic acceleration-time criterion based on the comparison between the stochastic acceleration time and the Hubble time, and (ii) a Fokker-Planck model for the isotropic ion distribution. This approach allows us to quantify both the redshift at which stochastic acceleration becomes cosmologically viable and the magnitude of any resulting nonthermal population. The central question we address is whether instability-enhanced scattering during cosmological magnetogenesis can reduce the stochastic acceleration time below the cosmological expansion timescale over any redshift interval prior to the widespread emergence of structure-formation shocks. In particular, we derive a critical condition for cosmologically viable second-order Fermi acceleration and follow the redshift evolution of the ion distribution under the combined effects of stochastic diffusion and adiabatic expansion. This framework enables a self-consistent assessment of (i) the epoch at which stochastic acceleration may become competitive with cosmic expansion, and (ii) the magnitude of any resulting suprathermal population. Energy losses due to Coulomb interactions are known to play an important role in shaping the low-energy spectrum of charged particles in diffuse astrophysical plasmas [e.g., 9, 30, 26]. In particular, Coulomb collisions can efficiently thermalize sub-relativistic particles and suppress the formation of nonthermal tails in weakly collisional environments. However, their impact on stochastic acceleration in the cosmological context—especially during the magnetogenesis phase prior to nonlinear structure formation—has not been systematically quantified. In this work, we incorporate Coulomb losses into both the analytic framework and the Fokker–Planck model, and examine how collisional cooling modifies particle injection and the resulting nonthermal population.
The paper is organized as follows. In Section 2, we derive the stochastic acceleration timescale within the instability-mediated magnetogenesis framework and obtain an analytic criterion for the cosmological viability of stochastic acceleration by considering adiabatic losses associated with cosmic expansion. This provides an optimistic upper bound on the efficiency of particle acceleration and defines a critical magnetic field and a corresponding turn-on redshift. In Section 3, we extend this analysis by including Coulomb energy losses in a fully ionized intergalactic medium. In Section 4, we solve a Fokker–Planck equation to quantify the resulting ion distribution. This allows us to assess the impact of collisional cooling on particle injection and to identify an effective energy threshold for stochastic acceleration. In Section 5, we compare the stochastic acceleration timescale with that of DSA in the Bohm-diffusion limit. Finally, Section 6 summarizes our results and discusses their cosmological implications.
2 Cosmological Viability of Stochastic Acceleration During Pressure-Anisotropy-Driven Magnetogenesis
Magnetic-field amplification mediated by pressure-anisotropy-driven instabilities can substantially modify particle transport properties in weakly magnetized plasmas. Since magnetic fields play a central role in regulating particle scattering, the growth of cosmic magnetic fields prior to the onset of large-scale structure formation may provide conditions for stochastic particle acceleration. In this section, we assess the cosmological viability of such acceleration by deriving a necessary condition set by the expansion of the universe. Specifically, we compare the stochastic acceleration timescale with the Hubble time, which defines the minimum requirement for particle energization on cosmological timescales. This approach provides an optimistic estimate of the acceleration efficiency and serves as a baseline against which additional physical effects, such as collisional energy losses, can be evaluated in subsequent sections. Throughout this work, we focus on ions, specifically protons, as the dominant baryonic component of the plasma, which sets the relevant mass and gyrofrequency scales for the acceleration process.
2.1 Acceleration Timescale
We adopt a stochastic acceleration scenario in which particles gain energy through repeated scattering off moving magnetic irregularities. In general, the acceleration timescale depends on the velocity fluctuations at the scale with which particles interact, i.e., at scales comparable to the particle mean free path $\lambda(E,z)$. The acceleration timescale can therefore be expressed schematically as
$t_{\rm acc}(E,z)\sim\left(\frac{c}{v_{\lambda}(z)}\right)^{2}\frac{\lambda(E,z)}{c},$ (2.1)
where $v_{\lambda}(z)$ denotes the characteristic turbulent velocity at the scale $\lambda$. In the present work, we adopt an optimistic approximation in which particles are assumed to sample velocity fluctuations comparable to the outer-scale turbulent velocity, $v_{\lambda}\sim v_{\rm tur}$. Under this assumption, the acceleration timescale reduces to
$t_{\rm acc}(E,z)\sim\left(\frac{c}{v_{\rm tur}(z)}\right)^{2}\frac{\lambda(E,z)}{c},$ (2.2)
which should be regarded as a lower-bound estimate of the true acceleration time. In realistic turbulence, where $v_{\lambda}<v_{\rm tur}$ for $\lambda$ below the outer scale, the actual acceleration time is expected to be longer, thereby further suppressing particle acceleration.
The mean free path is determined by the effective scattering frequency,
$\lambda(E,z)\approx\frac{v(E)}{\nu_{\rm eff}(z)}.$ (2.3)
The effective scattering rate is defined as
$\nu_{\rm eff}(z)\equiv\nu_{ii}(z)+\nu_{\rm scatt}(z),$ (2.4)
where $\nu_{ii}$ is the ion-ion (proton-proton) Coulomb collision frequency and $\nu_{\rm scatt}$ represents the scattering induced by pressure-anisotropy-driven instabilities. Following the instability-mediated magnetogenesis framework [e.g., 13], the scattering frequency in the regime where plasma instabilities dominate can be expressed as
$\nu_{\rm scatt}\sim\left(|\Delta|-2\beta^{-1}\right)^{3/2}\omega_{\rm ci},$ (2.5)
where $\Delta$ is the pressure anisotropy factor, $\beta$ is the plasma beta parameter, and $\omega_{\rm ci}=eB/m_{i}c$ is the ion gyrofrequency associated with the evolving magnetic field. This expression characterizes the pitch-angle scattering rate induced by pressure-anisotropy-driven microinstabilities, and is most directly applicable to particles with velocities comparable to the thermal speed. For higher-energy particles, the effective scattering rate is expected to depend on particle rigidity, leading to an energy-dependent diffusion coefficient. In the present work, we adopt Eq. (2.5) as an effective scattering rate that sets the normalization of stochastic acceleration. This corresponds to an optimistic assumption that maximizes the scattering efficiency. Any additional energy dependence that reduces the scattering rate at high energies would increase the acceleration time, thereby further suppressing particle acceleration.
Substituting Eq. (2.3) into Eq. (2.2), the acceleration timescale becomes
$t_{\rm acc}(z)\approx\left(\frac{c}{v_{\rm tur}(z)}\right)^{2}\frac{v(E)}{c\nu_{\rm eff}(z)}.$ (2.6)
This expression highlights that magnetic-field amplification, through its impact on $\omega_{\rm ci}$ and $\nu_{\rm scatt}$, can significantly shorten the acceleration time. Throughout this work, we adopt the scattering-dominated regime, consistent with the transition discussed in previous studies [6, 13]. In the very early stages of magnetic-field amplification, collisional effects can in principle influence the plasma dynamics. However, as the magnetic field grows and the ion gyrofrequency $\omega_{\rm ci}\propto B$ increases, the ratio $\nu_{ii}/\omega_{\rm ci}\propto(1+z)^{3}B^{-1}$ rapidly decreases. For redshifts $z\lesssim 10$, where the baryon density has already declined substantially and magnetic amplification has progressed, the condition $\nu_{ii}/\omega_{\rm ci}\ll 1$ is naturally satisfied. In this regime, pitch-angle scattering mediated by plasma instabilities dominates over Coulomb collisions ($\nu_{ii}\ll\nu_{\rm scatt}$). The effective scattering rate therefore satisfies $\nu_{\rm eff}\approx\nu_{\rm scatt}$, making the stochastic acceleration framework adopted here self-consistent.
To determine whether stochastic acceleration is viable before the formation of large-scale shocks, we compare the acceleration timescale with the Hubble time,
$t_{H}(z)=H^{-1}(z),$ (2.7)
where
$H(z)=H_{0}\sqrt{\Omega_{m}(1+z)^{3}+\Omega_{\Lambda}}.$ (2.8)
A necessary condition for efficient cosmological acceleration is therefore
$t_{\rm acc}(E,z)<t_{H}(z).$ (2.9)
Combining the above expressions yields the dimensionless ratio
$\frac{t_{\rm acc}}{t_{H}}\approx\left(\frac{c}{v_{\rm tur}}\right)^{2}\left(\frac{v(E)}{c}\right)\frac{H(z)}{\nu_{\rm eff}(z)}.$ (2.10)
When this ratio falls below unity, particles can be accelerated to relativistic energies within a cosmological expansion time. Conversely, if $t_{\rm acc}>t_{H}$, the expansion of the universe limits the achievable particle energies. From Eq. (2.10), one finds that the acceleration rate is related to the effective scattering rate as
$t_{\rm acc}^{-1}\sim\left(\frac{v_{\rm tur}}{c}\right)^{2}\left(\frac{c}{v}\right)\nu_{\rm eff},$ (2.11)
which reduces to $t_{\rm acc}^{-1}\sim(v_{\rm tur}/c)^{2}\nu_{\rm eff}$ in the relativistic limit ($v\sim c$). This shows that even in the presence of strong scattering, the acceleration efficiency is suppressed by the small ratio $(v_{\rm tur}/c)^{2}$. As a result, rapid pitch-angle scattering does not necessarily imply efficient particle acceleration.
Equation (2.9) provides a necessary condition for stochastic acceleration on cosmological timescales. However, it is not sufficient, as additional energy loss processes can impose more stringent constraints, particularly at low energies. In the low-energy regime, the characteristic timescales exhibit a clear hierarchy. Pitch-angle scattering is dominated by instability-driven processes, satisfying $\nu_{ii}\ll\nu_{\rm scatt}$, while Coulomb interactions can still strongly affect the energy evolution of particles. In particular, the Coulomb energy loss rate can be comparable to or exceed the stochastic acceleration rate, $t_{C}^{-1}\gtrsim t_{\rm acc}^{-1}$, thereby suppressing efficient energization of sub-relativistic particles. A more general requirement for particle acceleration is therefore
$t_{\rm acc}(E,z)<\min\left[t_{H}(z),t_{C}(E,z)\right],$ (2.12)
where $t_{C}\equiv E/|\dot{E}_{C}|$ is the characteristic Coulomb energy loss timescale. This condition highlights that, even in the presence of efficient pitch-angle scattering, stochastic acceleration can be strongly suppressed at low energies due to collisional cooling. In this regime, particles undergo frequent scattering but gain energy only slowly, while Coulomb losses efficiently remove energy, preventing their injection into the nonthermal population. In the present analysis, we first use the comparison with the Hubble time to identify the cosmological regime where acceleration may become viable. The impact of Coulomb losses is incorporated in Section 3, where the acceleration condition is refined to include energy-dependent cooling effects that introduce an effective injection threshold for stochastic acceleration.
2.2 Implications for Pre-Shock CR Production
Eqs. (2.2)-(2.10) suggest that instability-driven scattering during magnetogenesis may create a channel for CR production prior to structure formation shocks. As the magnetic field grows, the associated increase in gyrofrequency enhances the scattering rate, thereby reducing the acceleration timescale. This feedback implies that even moderately amplified magnetic fields could enable stochastic acceleration in the early universe.
In this framework, the key quantity controlling the onset of acceleration is the ratio $H(z)/\nu_{\rm eff}(z)$. If instability-mediated scattering becomes sufficiently strong, stochastic acceleration may operate well before the emergence of collisionless shocks, potentially establishing a seed population of CRs that later participate in DSA. The results presented here therefore motivate a quantitative exploration of particle acceleration during cosmological magnetogenesis and its implications for the origin of the earliest nonthermal particle populations. To evaluate the viability of stochastic acceleration during pressure-anisotropy-driven magnetogenesis, we adopt a minimal self-consistent closure of the background plasma parameters.
We parametrize the turbulent velocity as a fixed fraction of the sound speed,
$v_{\rm tur}(z)=\epsilon\,c_{s}(z),$ (2.13)
where $\epsilon<1$ is a dimensionless parameter and
$c_{s}(z)=\sqrt{\frac{\gamma k_{B}T(z)}{\mu m_{p}}}$ (2.14)
is the adiabatic sound speed of the background plasma. We adopt $\mu\approx 0.59$ appropriate for a fully ionized primordial plasma. Observations and simulations of diffuse astrophysical plasmas, such as the intracluster medium, suggest that gas motions are typically subsonic relative to the sound speed. For example, the Hitomi satellite measured a velocity dispersion of $\sim 160,{\rm km,s^{-1}}$ in the Perseus cluster core, indicating subsonic turbulence [14], while numerical studies likewise find turbulent velocities that are a fraction of the sound speed [8]. Although these results pertain to gas that has been processed by structure-formation shocks in the low-redshift Universe, they provide a useful reference for the characteristic magnitude of turbulent motions in weakly magnetized plasmas. Motivated by this, we adopt a representative range $\epsilon=0.05$$0.5$ in the present study.
We assume that the pressure anisotropy remains near a quasi-stable value
$\Delta(z)\approx\Delta_{0}=\text{const}.$ (2.15)
Near the marginal stability condition, the plasma beta satisfies
$\beta(z)\sim\frac{2}{|\Delta_{0}|},$ (2.16)
which provides a closure relation between $\beta$ and the anisotropy amplitude. Taking into account the system for particle acceleration, the anisotropy factor of the system $\Delta=\Delta_{0}+\delta\Delta$ satisfies $|\Delta|>2\beta^{-1}$, where $\delta\Delta\ll\Delta_{0}$ represents a perturbation around the quasi-stable value $\Delta_{0}$, which is treated as a free parameter that drives the instability. The supercritical excess above the marginal threshold $|\Delta_{0}|\sim 2\beta^{-1}$ is
$\Delta_{\rm ex}(\delta,\beta)\equiv|\Delta_{0}+\delta\Delta|-|\Delta_{0}|=|\Delta_{0}+\delta\Delta|-2\beta^{-1},$ (2.17)
which reduces to $\Delta_{\rm ex}=\delta\Delta$ for $\Delta_{0}>0$ and $\delta\Delta\geq 0$. In high-$\beta$ astrophysical plasmas such as the intracluster medium (ICM), where $\beta\sim\mathcal{O}(10^{2})$ [e.g., 32, 33, 34, 28, 18, 19], this corresponds to a typical anisotropy level $\Delta_{0}\sim\mathcal{O}(10^{-2})$.
The magnetic field evolves according to the instability-driven growth equation derived in the scattering-dominated regime [13],
$\frac{dB}{dt}$ $\approx\omega_{\rm ci,0}\frac{\left(|\Delta|-2\beta^{-1}\right)^{3/2}}{1+|\Delta|}\left(\frac{B}{B_{0}}\right)^{2}$
$+\omega_{\rm ci,0}\frac{|\delta\Delta|^{3/2}}{1+|\Delta|}e^{-\Gamma_{d}t-\frac{1}{2}\Gamma_{d}/\Gamma_{c}^{(\rm inst)}}\left(\frac{B}{B_{0}}\right)^{2},$ (2.18)
where the second term represents the contribution from perturbations of the pressure anisotropy. Here $\Gamma_{d}$ denotes the effective damping rate of the anisotropy due to pitch-angle scattering by plasma instabilities, whereas $\Gamma_{c}^{(\rm inst)}$ represents the characteristic rate at which large-scale plasma motions generate pressure anisotropy. $B_{0}$ is the seed magnetic field and $\omega_{\rm ci,0}$ corresponds to the ion gyrofrequency evaluated at $B_{0}$.
In high-$\beta$ weakly magnetized plasmas, pressure anisotropy is expected to be regulated near the instability threshold by rapid pitch-angle scattering, so that the strong-damping limit ($\Gamma_{d}\gg\Gamma_{c}^{(\rm inst)}$) provides an appropriate closure. For small perturbations of the anisotropy ($|\delta\Delta/\Delta|\ll 1$), the perturbation-driven contribution becomes negligible. While the detailed magnetic-field growth timescale can depend on the relative magnitude of the damping and perturbation terms [13], these effects primarily modify the growth rate without significantly altering the overall magnetic-field evolution. The magnetic-field evolution therefore reduces to
$\frac{dB}{dt}\approx\omega_{\rm ci,0}\frac{\left(|\Delta|-2\beta^{-1}\right)^{3/2}}{1+|\Delta|}\left(\frac{B}{B_{0}}\right)^{2}.$ (2.19)
This nonlinear growth implies that once the instability condition $|\Delta|>2\beta^{-1}$ is satisfied, the magnetic field amplifies superlinearly with $B$, leading to rapid amplification during the magnetogenesis phase.
2.3 Analytic constraint on the critical magnetic field
A necessary condition for efficient cosmological acceleration is that the acceleration time be shorter than the Hubble time, $t_{\rm acc}<t_{H}=H^{-1}(z)$. Combining Eqs.  (2.5), (2.10), and (2.17) yields an analytic lower bound on the magnetic field:
$B>B_{\rm crit}(z)\approx\frac{m_{i}c}{e}\left(\frac{c}{v_{\rm tur}(z)}\right)^{2}\left(\frac{v(E)}{c}\right)\frac{H(z)}{\Delta_{\rm ex}^{3/2}}.$ (2.20)
This condition defines a critical magnetic field required for stochastic acceleration to compete with cosmological expansion, and therefore represents a necessary (but not sufficient) criterion for particle acceleration. Eq. (2.20) makes explicit the strong parametric dependence $B_{\rm crit}\propto H(z)\,\epsilon^{-2}c_{s}^{-2}(z)\Delta_{\rm ex}^{-3/2}$: unless the turbulent motions are sufficiently strong and/or the anisotropy is strongly supercritical, the magnetic field required for $t_{\rm acc}<t_{H}$ becomes unrealistically large, implying that pressure-anisotropy-driven instabilities alone are unlikely to produce a significant pre-shock CR population under generic cosmological conditions.
Using Eq. (2.20), we can derive an explicit redshift dependence of the critical field required for $t_{\rm acc}<t_{H}$. At $z\gtrsim 1$ (matter-dominated era) we approximate
$H(z)\simeq H_{0}\sqrt{\Omega_{m}}\,(1+z)^{3/2}.$ (2.21)
Adopting a parametrized thermal history,
$T(z)=T_{0}\left(\frac{1+z}{1+z_{0}}\right)^{\alpha},$ (2.22)
the sound speed satisfies $c_{s}^{2}(z)\propto T(z)\propto(1+z)^{\alpha}$. Eq. (2.20) then yields
$B_{\rm crit}(z)\propto\frac{H(z)}{c_{s}^{2}(z)}\propto(1+z)^{\frac{3}{2}-\alpha},$ (2.23)
or equivalently
$B_{\rm crit}(z)=B_{\rm crit}(z_{0})\left(\frac{1+z}{1+z_{0}}\right)^{\frac{3}{2}-\alpha},$ (2.24)
where $B_{\rm crit}(z_{0})$ collects the normalization factors, including $\epsilon^{-2}$ and $\Delta_{\rm ex}^{-3/2}$. The parameter $\alpha$ characterizes the thermal history of the background plasma through $T(z)\propto(1+z)^{\alpha}$. In the absence of heating processes, adiabatic expansion of non-relativistic particles leads to $T\propto a^{-2}$, corresponding to $\alpha=2$. However, in realistic cosmological environments, additional heating mechanisms—such as photoheating, residual Compton heating, and structure formation—can partially offset adiabatic cooling, resulting in a shallower temperature evolution with $\alpha<2$. For this reason, it is reasonable to consider a range of thermal histories with $1\lesssim\alpha\lesssim 2$. In the illustrative case $\alpha=1$, corresponding to a moderately heated medium, the critical field scales as $B_{\rm crit}(z)\propto(1+z)^{1/2}$, implying a slowly increasing threshold toward higher redshift. By contrast, for purely adiabatic evolution ($\alpha=2$), one obtains $B_{\rm crit}(z)\propto(1+z)^{-1/2}$.
2.4 Definition of the CR turn-on redshift
A useful way to quantify when pressure-anisotropy-driven scattering can begin to enable cosmological stochastic acceleration is to define a ”CR turn-on redshift” $z_{\rm on}$. We define $z_{\rm on}$ as the epoch at which the stochastic acceleration time becomes comparable to the Hubble time,
$\left.\frac{t_{\rm acc}(z)}{t_{H}(z)}\right|_{z=z_{\rm on}}=1,\qquad t_{H}(z)=H^{-1}(z).$ (2.25)
This condition provides a necessary criterion for stochastic acceleration on cosmological timescales. For $z>z_{\rm on}$ the expansion time is too short ($t_{\rm acc}>t_{H}$) and stochastic acceleration is inefficient, while for $z<z_{\rm on}$ acceleration may in principle become viable. However, this condition is not sufficient, as additional energy loss processes—such as Coulomb and ionization losses—can impose more stringent constraints, particularly at low energies. Therefore, the $z_{\rm on}$ defined here should be interpreted as an upper bound on the redshift below which stochastic acceleration may operate.
Figure 1: Redshift evolution of the magnetic field $B(z)$ (black) compared with the analytic critical field $B_{\rm crit}(z)$ for two different thermal histories: $\alpha=1$ (solid red) and $\alpha=2$ (dashed red), corresponding to a moderately heated medium and purely adiabatic evolution, respectively. Their intersection defines the CR turn-on redshift $z_{\rm on}$. The curves are shown for fiducial parameters $B_{0}=10^{-18}$ G, $\delta\Delta/\Delta_{0}=0.1$, $\Delta_{0}\sim 2\beta^{-1}=10^{-2}$, and $\epsilon=0.2$. Here, the critical field is evaluated in the relativistic limit ($v\sim c$), providing a reference estimate for particle acceleration. The resulting $z_{\rm on}$ depends only weakly on the assumed thermal history.
Figure 2: Dependence of the CR turn-on redshift $z_{\rm on}$ on (a) the turbulence amplitude $\epsilon$ and (b) the pressure-anisotropy perturbation $\delta\Delta$. All other parameters are fixed to the fiducial values used in Figure 1. Over the explored parameter range, $z_{\rm on}$ varies only weakly, remaining close to $z_{\rm on}\sim 1.7$.
Using the analytic constraint derived above, the turn-on condition can be written equivalently as a crossing between the evolving magnetic field and the critical field:
$B(z_{\rm on})=B_{\rm crit}(z_{\rm on}),$ (2.26)
where $B_{\rm crit}(z)$ is defined by Eq. (2.20). In the matter-dominated regime $H(z)\propto(1+z)^{3/2}$ and for a parametrized thermal history $T(z)\propto(1+z)^{\alpha}$ (so that $c_{s}^{2}\propto(1+z)^{\alpha}$), one finds $B_{\rm crit}(z)\propto(1+z)^{3/2-\alpha}$. For illustrative purposes, if the magnetic-field evolution can be approximated by $B(z)=B_{0}(1+z)^{p}$, the turn-on redshift follows from Eq. (2.26) as
$1+z_{\rm on}=\left(\frac{C}{B_{0}}\right)^{\!\frac{1}{\,p-\left(\frac{3}{2}-\alpha\right)}},$ (2.27)
where $C$ collects the normalization factors in $B_{\rm crit}$, including the strong dependences $C\propto\epsilon^{-2}\Delta_{\rm ex}^{-3/2}$. A finite $z_{\rm on}$ therefore exists only if the magnetic field grows sufficiently rapidly with redshift, $p>(3/2-\alpha)$; otherwise the system remains in the acceleration-inefficient regime at all epochs.
To illustrate this behavior explicitly, we compute the magnetic-field evolution $B(z)$ using the unperturbed scattering-dominated solution from our previous work and compare it with the analytic critical field $B_{\rm crit}(z)$. Here, the critical field is evaluated in the relativistic limit ($v\sim c$), providing a baseline estimate for particle acceleration. Figure 1 shows the resulting redshift evolution for a representative set of parameters, including two different thermal histories characterized by $\alpha=1$ and $\alpha=2$. At high redshift, the magnetic field remains far below the critical value in both cases, implying $t_{\rm acc}\gg t_{H}$ and therefore inefficient stochastic acceleration. As cosmic time progresses, nonlinear magnetic amplification driven by pressure-anisotropy instabilities leads to a rapid growth of $B(z)$. The CR turn-on redshift $z_{\rm on}$ is defined by the crossing $B(z)=B_{\rm crit}(z)$, marking the epoch at which stochastic acceleration becomes viable on cosmological timescales. In all cases considered here, the crossing occurs at $z_{\rm on}\sim 1.7$, corresponding to a magnetic field strength of order $10^{-9}$ G. Notably, the location of this crossing depends only weakly on the assumed thermal history. This can be understood from the different redshift scalings of the two quantities: while $B_{\rm crit}(z)$ evolves only as a power law in $(1+z)$, the magnetic field undergoes rapid nonlinear amplification. As a result, the onset of stochastic acceleration is primarily governed by instability-driven magnetization, with the thermal history playing only a secondary role. Importantly, this epoch coincides with the onset of nonlinear structure formation, suggesting that efficient cosmic-ray production is naturally associated with structure-formation shocks rather than with earlier microinstability-driven stochastic processes.
Figure 2 illustrates the dependence of the turn-on redshift on the turbulence amplitude $\epsilon$ and on the pressure anisotropy, parameterized by $\delta\Delta/\Delta_{0}$ and $\Delta_{0}\sim 2\beta^{-1}$. Despite the explicit appearance of both quantities in the analytic expression for $B_{\rm crit}(z)$, the resulting variation in $z_{\rm on}$ is remarkably small over the considered parameter range. In particular, even when $\epsilon$ and $\delta\Delta$ are varied by factors of a few, the turn-on redshift remains confined to a narrow interval around $z_{\rm on}\sim 1.7$. This behavior reflects the fact that $B(z)$ exhibits a rapid nonlinear growth phase, so that modest vertical shifts in $B_{\rm crit}(z)$ translate into only small horizontal shifts in the crossing redshift. The onset of cosmologically viable stochastic acceleration is therefore robust against moderate changes in the underlying microphysical parameters. Importantly, the inferred $z_{\rm on}$ remains close to the epoch of nonlinear structure formation, reinforcing the conclusion that efficient CR production is naturally associated with structure-formation shocks rather than with earlier microinstability-driven processes.
2.5 Maximum attainable energy during cosmological magnetogenesis
While the condition $t_{\rm acc}(z)<t_{H}(z)$ determines when stochastic acceleration becomes cosmologically viable, it does not specify the maximum particle energy that can be reached. To quantify the physical significance of the turn-on epoch, we estimate the maximum ion energy attainable before the expansion of the Universe limits further acceleration.
To estimate the largest particle energies that could in principle be reached, we adopt an optimistic upper bound for the particle mean free path by assuming $\lambda\sim r_{L}$, where $r_{L}$ is the Larmor radius. This corresponds to the strong-scattering limit $\nu_{\rm scatt}\sim\omega_{\rm ci}$, which can arise when instability-driven pitch-angle scattering dominates. In this limit, the spatial diffusion coefficient becomes
$D(E)\sim\frac{1}{3}r_{L}(E)c,\qquad r_{L}(E)=\frac{pc}{eB}\approx\frac{E}{eB},$ (2.28)
for relativistic ions with $E\approx pc$. In this limit, the second-order Fermi acceleration time can be expressed as
$t_{\rm acc,2}(E,z)\approx\left(\frac{c}{v_{\rm tur}(z)}\right)^{2}\frac{r_{L}(E)}{c}=\left(\frac{c}{v_{\rm tur}(z)}\right)^{2}\frac{E}{eB(z)c}.$ (2.29)
This expression captures the essential energy dependence: higher-energy particles require longer acceleration times.
Figure 3: Redshift evolution of the maximum attainable ion energy $E_{\max}(z)$ computed from Eq. (2.31) for turbulence amplitudes $\epsilon=0.05$ (black), $0.1$ (red), and $0.2$ (blue). Near the turn-on epoch $z_{\rm on}\sim 1.7$, $E_{\max}$ reaches $\sim 10^{2}\,\mathrm{GeV}$.
The maximum attainable energy at a given redshift is obtained by requiring that the acceleration time not exceed the Hubble time,
$t_{\rm acc,2}(E_{\max},z)=t_{H}(z)=H^{-1}(z).$ (2.30)
Substituting the parametrization $v_{\rm tur}(z)=\epsilon c_{s}(z)$ into Eq. (2.29), the maximum attainable energy becomes
$E_{\max}(z)\approx eB(z)c\epsilon^{2}\left(\frac{c_{s}(z)}{c}\right)^{2}H^{-1}(z).$ (2.31)
In the matter-dominated regime, $H(z)\propto(1+z)^{3/2}$. Adopting the parametrized thermal history $T(z)\propto(1+z)^{\alpha}$, so that $c_{s}^{2}(z)\propto(1+z)^{\alpha}$, we find
$E_{\max}(z)\propto B(z)\,(1+z)^{\alpha-3/2}.$ (2.32)
Thus, even if the magnetic field grows rapidly, cosmological expansion suppresses the maximum achievable energy at high redshift.
Figure 3 shows the redshift evolution of the maximum attainable ion energy for representative values of the turbulence parameter $\epsilon$. Evaluated near the turn-on epoch $z_{\rm on}\sim 1.7$, where $B(z_{\rm on})\sim 10^{-9}\,\mathrm{G}$ in our fiducial model, Eq. (2.31) implies that ion energies can, under optimistic optimistic assumptions corresponding to the strong-scattering limit, reach values up to $\sim 10^{2}\,\mathrm{GeV}$. The strong dependence $E_{\max}\propto\epsilon^{2}$ implies that larger turbulence amplitudes lead to higher maximum energies, while the overall redshift dependence remains unchanged. At higher redshift ($z\gtrsim z_{\rm on}$), the magnetic field remains weak and the Hubble time short, so that $E_{\max}$ is strongly suppressed and remains far below the relativistic regime. Thus, even under optimistic assumptions, cosmological stochastic acceleration cannot generate a substantial high-energy particle population prior to the onset of nonlinear structure formation. We emphasize that Eq. (2.31) provides an upper bound. Additional physical effects—including Coulomb and ionization losses, incomplete turbulence development, or deviations from Bohm-like diffusion—would further reduce the achievable energy. Consequently, while stochastic acceleration may in principle produce $\mathcal{O}(10$-$10^{2})\,\mathrm{GeV}$ ions near $z_{\rm on}\sim 1.7$, the emergence of structure-formation shocks remains a far more efficient and robust mechanism for accelerating CRs to higher energies.
3 Refined Constraints Including Coulomb Losses in a Fully Ionized Medium
In Section 2, we derived the conditions for stochastic acceleration by considering adiabatic losses associated with cosmological expansion. This provides an optimistic estimate of the acceleration efficiency, corresponding to a lower bound on the acceleration time and a minimal requirement for particle energization. In this section, we refine the analysis by including Coulomb losses, which arise from interactions with background electrons and become important at sub-relativistic energies. We emphasize that this effect is distinct from the role of Coulomb collisions in pitch-angle scattering: while Coulomb scattering is negligible compared to instability-driven scattering (i.e., $\nu_{ii}\ll\nu_{\rm scatt}$), Coulomb interactions can still dominate the energy evolution of particles through collisional energy losses, in the sense that the Coulomb loss rate can exceed the stochastic acceleration rate ($t_{C}^{-1}\gtrsim t_{\rm acc}^{-1}$). We focus on a fully ionized medium, appropriate for the post-reionization intergalactic medium. Since the turn-on redshift derived in this work lies at $z_{\rm on}\sim 1.7$, well below the epoch of reionization ($z\sim 6$), this assumption is well justified. In such an environment, Coulomb interactions dominate the collisional energy losses, allowing us to isolate the impact of collisional cooling on stochastic acceleration in a physically well-defined and tractable regime. We note that additional ionization losses in partially neutral regions would further shorten the cooling timescale and thus strengthen the constraints derived below.
3.1 Energy loss processes in the pre-structure intergalactic medium
In addition to adiabatic losses due to cosmological expansion, charged particles in the pre-structure intergalactic medium experience energy losses through Coulomb interactions with free electrons and, in partially neutral regions, through ionization processes. The characteristic energy loss timescales are defined as
$t_{\rm C}(E,z)\equiv\frac{E}{|\dot{E}_{\rm C}(E,z)|},\qquad t_{\rm ioniz}(E,z)\equiv\frac{E}{|\dot{E}_{\rm ioniz}(E,z)|}.$ (3.1)
Both processes are most effective at sub-relativistic energies and scale inversely with the ambient density, $t_{\rm C},t_{\rm ioniz}\propto n^{-1}(z)$. Since the baryon density evolves as $n(z)\propto(1+z)^{3}$, the corresponding cooling times decrease rapidly toward higher redshift. In ionized regions of the intergalactic medium, Coulomb interactions with free electrons dominate the collisional energy losses, while ionization losses become important only in partially neutral environments. In the following, we adopt a fully ionized medium as a fiducial model, appropriate for the post-reionization universe, and therefore focus on Coulomb losses as the primary non-adiabatic cooling process. Coulomb interactions of ions are generally dominated by scattering off thermal electrons [e.g., 35], owing to the cumulative effect of frequent small-angle encounters with light particles. Accordingly, we approximate the Coulomb energy loss rate using the electron contribution.
For sub-relativistic and mildly relativistic ions in an ionized plasma, the Coulomb energy loss rate is dominated by interactions with background electrons and can be approximated as
$\dot{E}_{\rm C}(E,z)\approx\frac{4\pi e^{4}n_{e}(z)\ln\Lambda}{m_{e}v(E)},$ (3.2)
where $n_{e}(z)$ is the electron number density, $\ln\Lambda$ is the Coulomb logarithm, and $v(E)$ is the particle velocity. This leads to a characteristic Coulomb loss timescale
$t_{\rm C}(E,z)\approx\frac{Em_{e}v(E)}{4\pi e^{4}n_{e}(z)\ln\Lambda}.$ (3.3)
This expression explicitly shows that Coulomb losses become increasingly efficient at low energies due to the inverse dependence on particle velocity.
The total effective loss rate is given by
$t_{\rm loss}^{-1}(E,z)=t_{H}^{-1}(z)+t_{C}^{-1}(E,z),$ (3.4)
so that in the limits $t_{C}\ll t_{H}$ or $t_{H}\ll t_{C}$, the effective loss timescale reduces to the shorter of the two. We note that the inclusion of ionization losses in partially neutral regions would further shorten the cooling timescale at low energies, thereby strengthening the constraints derived below.
3.2 Refined critical magnetic field
Including these additional loss processes, the condition for stochastic acceleration becomes
$t_{\rm acc}(E,z)<t_{\rm loss}(E,z).$ (3.5)
This leads to a refined critical magnetic field, $B_{\rm crit}^{\rm eff}(E,z)$, defined implicitly by
$t_{\rm acc}(E,z)=t_{\rm loss}(E,z).$ (3.6)
Using the optimistic acceleration timescale adopted in Section 2, we obtain an explicit expression for the refined critical field,
$B_{\rm crit}^{\rm eff}(E,z)\approx\frac{m_{i}c}{e}\left(\frac{c}{v_{\rm tur}(z)}\right)^{2}\left(\frac{v(E)}{c}\right)\frac{1}{\Delta_{\rm ex}^{3/2}t_{\rm loss}(E,z)}.$ (3.7)
This expression reduces to the Hubble-limited critical field when $t_{\rm loss}=t_{H}$,
$B_{\rm crit}(z)\approx\frac{m_{i}c}{e}\left(\frac{c}{v_{\rm tur}(z)}\right)^{2}\left(\frac{v(E)}{c}\right)\frac{H(z)}{\Delta_{\rm ex}^{3/2}}.$ (3.8)
More generally, the two are related by
$B_{\rm crit}^{\rm eff}(E,z)=\frac{B_{\rm crit}(z)}{H(z)t_{\rm loss}(E,z)}.$ (3.9)
Figure 4: Redshift evolution of the magnetic field $B(z)$ (black) compared with the refined critical field $B_{\rm crit}^{\rm eff}(E,z)$ for different particle energies including Coulomb losses. The gray curve shows the Hubble-limited result. Coulomb losses strongly suppress acceleration at low energies, while at high energies $B_{\rm crit}^{\rm eff}$ converges to the Hubble-limited case, indicating an effective energy threshold for stochastic acceleration.
Since Coulomb losses reduce the effective loss timescale at low energies, $t_{\rm loss}(E,z)\leq t_{H}(z)$, the refined critical field satisfies
$B_{\rm crit}^{\rm eff}(E,z)\gtrsim B_{\rm crit}(z),$ (3.10)
indicating that a stronger magnetic field is required for efficient acceleration when realistic energy losses are taken into account. In particular, in the regime where Coulomb losses dominate ($t_{\rm loss}\approx t_{\rm C}$), the critical field becomes explicitly energy dependent,
$B_{\rm crit}^{\rm eff}(E,z)\propto\frac{1}{v_{\rm tur}^{2}(z)}\frac{v(E)}{t_{\rm C}(E,z)\Delta_{\rm ex}^{3/2}},$ (3.11)
reflecting the fact that low-energy particles are more strongly affected by collisional cooling. This introduces an additional suppression of stochastic acceleration compared to the adiabatic-only estimate.
Figure 4 shows the redshift evolution of the refined critical magnetic field $B_{\rm crit}^{\rm eff}(E,z)$ for a range of particle energies, together with the magnetic-field evolution $B(z)$ and the Hubble-limited critical field. At low energies, Coulomb losses significantly reduce the effective loss timescale ($t_{\rm loss}\approx t_{\rm C}$), leading to a substantial increase in $B_{\rm crit}^{\rm eff}$. As a result, the required magnetic field for efficient stochastic acceleration becomes much larger than in the adiabatic-only case, and the crossing with $B(z)$ is either shifted to much lower redshift or entirely suppressed. This demonstrates that collisional cooling strongly inhibits particle acceleration near the injection scale. In contrast, at sufficiently high energies where $t_{\rm C}\gg t_{H}$, the effective loss timescale is governed by cosmological expansion ($t_{\rm loss}\approx t_{H}$), and $B_{\rm crit}^{\rm eff}$ converges to the Hubble-limited critical field derived in Section II. In this regime, stochastic acceleration becomes largely insensitive to Coulomb losses, and the corresponding turn-on redshift approaches the relativistic baseline shown in Figure 1. These results clearly demonstrate that Coulomb losses introduce a strong energy dependence in the conditions for stochastic acceleration, primarily affecting low-energy particles while leaving high-energy particles largely unaffected.
The comparison between $B(z)$ and $B_{\rm crit}^{\rm eff}(E,z)$ further implies the existence of an effective low-energy threshold for stochastic acceleration. Below this threshold, Coulomb losses dominate over acceleration, preventing particles from gaining energy efficiently and suppressing their injection into the acceleration process. In the fiducial model considered here, this threshold is found to lie at approximately the ${\mathcal{O}}(10)$ keV level. Particles with energies below this scale experience rapid collisional cooling, such that $t_{\rm C}\ll t_{\rm acc}$ over the relevant redshift range. As a result, stochastic acceleration is effectively quenched. Conversely, for particles with energies above this threshold, the Coulomb loss timescale increases rapidly, allowing stochastic acceleration to operate and eventually approach the Hubble-limited regime. We therefore conclude that Coulomb losses impose an effective injection energy for stochastic acceleration, transforming the acceleration condition from a purely cosmological constraint into an energy-dependent threshold problem. At sufficiently high energies, however, where $t_{\rm C}\gg t_{H}$, the acceleration process becomes limited by cosmological expansion, and the maximum attainable energy remains essentially unchanged from the adiabatic-only case.
4 Pre-Structure Stochastic Acceleration in the Early Universe
To quantify the magnitude of any seed nonthermal population produced by the instability-assisted stochastic process discussed above, we introduce a transport model for the isotropic ion distribution function. The goal is not to construct a fully self-consistent cosmological CR population synthesis, but rather to translate the acceleration-time constraints into an explicit estimate of the resulting momentum-space spectrum and energy content.
4.1 Isotropic Fokker-Planck equation in an expanding background
Let $f(p,z)$ be the isotropic phase-space distribution of ions, normalized such that the differential number density is $dn=4\pi p^{2}f(p,z)\,dp$. Neglecting spatial gradients and assuming isotropic pitch-angle scattering, the momentum-space Fokker-Planck equation reads
$\frac{\partial f}{\partial t}$ $=\frac{1}{p^{2}}\frac{\partial}{\partial p}\left[p^{2}D_{pp}(p,z)\,\frac{\partial f}{\partial p}\right]$
$-\frac{1}{p^{2}}\frac{\partial}{\partial p}\left[p^{2}\,\dot{p}_{\rm loss}(p,z)\,f\right]+Q(p,z)-\frac{f}{t_{\rm esc}(p,z)},\,$ (4.1)
where $D_{pp}$ is the momentum diffusion coefficient associated with second-order Fermi acceleration, $\dot{p}_{\rm loss}$ represents systematic momentum losses, $Q$ is a source term (e.g., injection from the thermal pool), and $t_{\rm esc}$ is an effective escape time from the acceleration region. Cosmological expansion is incorporated by transforming from $t$ to redshift $z$ using
$\frac{dt}{dz}=-\frac{1}{(1+z)H(z)}\,$, (4.2)
with $H(z)=H_{0}\sqrt{\Omega_{\rm m}(1+z)^{3}+\Omega_{\Lambda}}$. Eq. (4.1) then becomes
$-(1+z)H(z)\,\frac{\partial f}{\partial z}$ $=\frac{1}{p^{2}}\frac{\partial}{\partial p}\left[p^{2}D_{pp}(p,z)\,\frac{\partial f}{\partial p}\right]$
$-\frac{1}{p^{2}}\frac{\partial}{\partial p}\left[p^{2}\,\dot{p}_{\rm loss}(p,z)\,f\right]+Q(p,z)\,.$ (4.3)
In a second-order Fermi process associated with the pressure-anisotropy-driven instabilities, which generate small-scale magnetic fluctuations that scatter particles, the characteristic acceleration time can be expressed in terms of $D_{pp}$ as
$D_{pp}(p,z)=\frac{p^{2}}{t_{\rm acc}(z)}~(p\gtrsim p_{\rm inj}),$ (4.4)
corresponding to a momentum diffusion coefficient proportional to $p^{2}$, with a redshift-dependent normalization determined by the microphysical scattering rate. Here, $p_{\rm inj}$ denotes the injection momentum for stochastic acceleration. This form represents the minimal choice consistent with the timescale analysis presented in Section 2. In the present model, we assume that the escape timescale is much longer than the characteristic acceleration and loss timescales, $t_{\rm esc}\gg t_{\rm acc},~t_{\rm loss}$. The characteristic loss timescale is defined as $t_{\rm loss}\equiv p/\dot{p}_{\rm loss}$. Cosmological expansion inevitably induces adiabatic momentum losses,
$\dot{p}_{\rm ad}(p,z)\approx-H(z)p.$ (4.5)
In addition, particles experience Coulomb losses due to interactions with background electrons, which are particularly important at sub-relativistic energies. The total momentum loss rate is therefore given by
$\dot{p}_{\rm loss}(p,z)=\dot{p}_{\rm ad}(p,z)+\dot{p}_{\rm C}(p,z),$ (4.6)
where the Coulomb momentum loss rate can be approximated as
$\dot{p}_{\rm C}(p,z)\approx-\frac{4\pi e^{4}n_{e}(z)\ln\Lambda}{m_{e}v^{2}(p)}.$ (4.7)
Here $n_{e}(z)$ is the electron number density, $\ln\Lambda$ is the Coulomb logarithm, and $v(p)$ is the particle velocity. This expression shows that Coulomb losses increase rapidly toward low velocities, leading to efficient thermalization of particles near the injection scale. Including both contributions, Eq. (4.3) becomes
$-(1+z)H(z),\frac{\partial f}{\partial z}$ $\approx\frac{1}{p^{2}}\frac{\partial}{\partial p}\left[p^{2}\frac{p^{2}}{t_{\rm acc}(z)}\frac{\partial f}{\partial p}\right]$
$-\frac{1}{p^{2}}\frac{\partial}{\partial p}\left[p^{2}\left(\dot{p}_{\rm ad}(p,z)+\dot{p}_{\rm C}(p,z)\right)f\right]$
$+Q(p,z).$ (4.8)
The Coulomb term introduces an additional sink of particle momentum at low energies, which can suppress the formation of a suprathermal tail by efficiently thermalizing particles near the injection scale.
We note that adiabatic momentum losses due to cosmological expansion are applicable to particles residing in the unbound intergalactic medium considered in this work. In contrast, particles confined within gravitationally bound systems, such as galaxies or galaxy clusters, do not experience cosmological adiabatic cooling. Since our analysis focuses on the pre-structure phase prior to nonlinear collapse, the assumption $\dot{p}_{\rm ad}\approx-H(z)p$ remains valid throughout the redshift interval of interest.
4.2 Numerical implementation and early-universe heating
To assess the magnitude of stochastic heating prior to the magnetic-field turn-on epoch, we solve Eq. (4.8) numerically over the redshift interval $z=10\rightarrow z_{\rm on}$, where the magnetic field remains well below its rapid-amplification regime. In this early phase, the goal is to determine whether the instability-assisted second-order Fermi process can produce a significant seed nonthermal component over a Hubble time.
The initial condition at $z=10$ is taken to be a non-relativistic Maxwellian distribution,
$f_{\rm MW}(p,z=10)=\frac{n_{i}}{(2\pi m_{i}k_{B}T_{0})^{3/2}}\exp\left(-\frac{p^{2}}{2m_{i}k_{B}T_{0}}\right),$ (4.9)
where $n_{i}$ denotes the ion number density. We emphasize that this setup is not intended to represent the volume-averaged thermal state of the universe at $z=10$, but rather a simplified diffuse ionized medium used as a fiducial environment for assessing the competition between stochastic acceleration and cooling processes. Accordingly, $T_{0}$ should be interpreted as an effective model parameter. In the fiducial case, we adopt $k_{B}T_{0}=0.86~{\rm eV}$, corresponding to a moderately heated plasma. No explicit source term is included ($Q=0$), and escape is neglected, so that the evolution is governed solely by stochastic momentum diffusion and energy losses. The magnetic field $B(z)$ is mapped to redshift using the cosmological relation $dt/dz=-[(1+z)H(z)]^{-1}$. In the early regime ($z\gtrsim z_{\rm on}$), $B(z)$ evolves smoothly and remains far below the explosive growth associated with $z<z_{\rm on}$, ensuring that $t_{\rm acc}(z)$ varies gradually.
Eq. (4.8) is solved on a logarithmic momentum grid using a fully implicit scheme in redshift. The equation is written in conservative flux form in momentum space,
$\mathcal{R}(f)=\frac{1}{p^{2}}\frac{\partial}{\partial p}\left[p^{2}D_{pp}\frac{\partial f}{\partial p}-p^{2}(\dot{p}_{\rm ad}+\dot{p}_{\rm C})f\right],$ (4.10)
and discretized such that particle number conservation is maintained up to numerical accuracy. The redshift step $\Delta z$ is mapped to an effective time step $\Delta t=-\Delta z/[(1+z)H(z)]$, ensuring consistency with the cosmological evolution.
Figure 5: Ion spectra $p^{4}f(p,z)$ from the Fokker–Planck model at $z=10,5,2,$ and $1.7$. Panel (a): full solution including Coulomb and adiabatic losses; panel (b): adiabatic-loss-only case. The initial distribution at $z=10$ is a Maxwellian with $k_{B}T_{0}=0.86~{\rm eV}$. The turbulent velocity is parameterized as $v_{\rm tur}=\epsilon c_{s}$ with $\epsilon=0.2$. Coulomb losses suppress the formation of a suprathermal tail by efficiently cooling low-energy particles.
Figure 5 shows the evolution of the particle spectra $p^{4}f(p,z)$ at $z=10,5,2,$ and $1.7$, comparing the full solution including Coulomb losses (left panel) with the adiabatic-loss-only case (right panel). In both cases, the thermal peak shifts toward lower momentum as redshift decreases, reflecting adiabatic cooling for which $p\propto a^{-1}\propto(1+z)$. The corresponding decrease in peak amplitude is likewise consistent with the dilution of the thermal energy density in an expanding background. In the absence of Coulomb losses (right panel), stochastic momentum diffusion produces a modest broadening of the high-momentum tail relative to pure adiabatic evolution. For stronger turbulence, the suprathermal component becomes slightly more pronounced as the system approaches the magnetic-field turn-on epoch, indicating that stochastic acceleration begins to compete with adiabatic cooling. In contrast, when Coulomb losses are included (left panel), the low-energy population is efficiently thermalized, and the formation of a suprathermal tail is strongly suppressed over the entire redshift range considered. Even for relatively strong turbulence, the distribution remains close to a Maxwellian, with only minimal deviation at high momenta. This behavior reflects the fact that Coulomb cooling dominates over stochastic acceleration near the injection scale, preventing particles from being efficiently accelerated out of the thermal pool. As the system approaches the magnetic-field turn-on epoch ($z\sim z_{\rm on}\sim 1.7$), the acceleration timescale decreases due to magnetic-field amplification. However, the results indicate that Coulomb losses continue to limit the development of a significant nonthermal component in the pre-shock phase. We therefore conclude that stochastic acceleration alone is unable to generate a substantial suprathermal seed population prior to the onset of structure-formation shocks.
We do not extend the Fokker–Planck integration below $z_{\rm on}$, since in this regime large-scale structure formation shocks are expected to dominate particle acceleration. The present model is therefore designed to quantify the pre-shock stochastic heating phase and to assess the extent to which a seed nonthermal population can be established before shock-driven acceleration takes over.
4.3 Limited impact of a pre-existing suprathermal population
While the preceding analysis assumes a purely thermal initial distribution, a weak suprathermal particle population may arise from a variety of processes operating in weakly magnetized plasmas, such as plasma instabilities, turbulence, or residual nonthermal components inherited from earlier evolutionary stages. Even a small nonthermal tail could in principle act as a seed population for subsequent acceleration.
To assess the sensitivity of the stochastic heating process to such initial conditions, we consider an alternative initial distribution consisting of a Maxwellian core with a small suprathermal power-law tail,
$f(p,z=10)=f_{\rm MW}(p,z=10)+Ap^{-s}\exp(-p/p_{\rm cut}).$ (4.11)
The normalization of the tail is fixed by matching it to the Maxwellian distribution at a reference momentum $p_{0}$, such that
$A,p_{0}^{-s}=\eta_{\rm tail}f_{\rm MW}(p_{0},z=10),$ (4.12)
where $\eta_{\rm tail}\ll 1$ controls the relative amplitude of the suprathermal component. This condition ensures a smooth connection between the thermal and nonthermal populations at $p=p_{0}$. We then evolve this distribution using Eq. (4.8). The slope of the pre-existing tail is taken to be $s=4.2$, representative of typical nonthermal spectra in collisionless plasmas, and consistent with expectations from both stochastic acceleration and shock-related processes. An exponential cutoff $p_{\rm cut}$ limits the high-momentum extent of the seed population.
Figure 6: Same as Figure 5, but including a weak pre-existing suprathermal power-law tail ($s=4.2$) in the initial distribution.
The resulting spectra are shown in Figure 6. We find that the overall evolution from $z=10$ to $z=2$ is largely insensitive to the presence of a weak pre-existing suprathermal population. In both panels, the ion distribution shifts toward lower momenta as the universe expands, reflecting adiabatic momentum losses $\dot{p}_{\rm ad}=-H(z)p$, for which $p\propto a^{-1}\propto(1+z)$. However, when Coulomb losses are included, a clear departure from purely adiabatic evolution appears at low momenta. In particular, particles with $p/(m_{i}c)\lesssim 5\times 10^{-3}$, corresponding to kinetic energies of order $\sim 10~{\rm keV}$, are efficiently thermalized by Coulomb interactions. In this regime, the cooling timescale is shorter than the acceleration timescale, and stochastic diffusion is effectively suppressed. As a result, the pre-existing power-law tail does not survive at low energies, and the distribution remains close to a Maxwellian core throughout the evolution. At higher momenta ($p/(m_{i}c)\gtrsim 5\times 10^{-3}$), the Coulomb loss timescale increases rapidly, allowing stochastic acceleration to operate. Nevertheless, even in this regime, the resulting modification of the spectrum remains modest over the interval $z=10\rightarrow 2$, indicating that cosmological cooling continues to dominate the overall evolution.
As the system approaches the magnetic-field turn-on epoch ($z\sim z_{\rm on}\sim 1.7$), the reduction of the stochastic acceleration timescale allows the diffusion term to partially compete with losses. However, the results indicate that Coulomb cooling still limits the efficient re-acceleration of the seed population. The presence of a pre-existing tail therefore does not significantly enhance the formation of a nonthermal component in the pre-shock phase. Instead, the early evolution is controlled primarily by the competition between Coulomb thermalization and cosmological expansion, with stochastic acceleration becoming effective only above the effective threshold set by Coulomb losses.
5 Comparison with shock acceleration in the Bohm-diffusion limit
Even if stochastic acceleration becomes formally viable after the turn-on epoch $z_{\rm on}$, it is important to assess whether it can compete with acceleration in structure-formation shocks. To this end, we compare the characteristic acceleration timescale of our stochastic acceleration mechanism with the standard DSA timescale assuming Bohm-like diffusion.
For a non-relativistic shock of speed $u_{s}$ (in the upstream frame), the DSA acceleration time can be written as [e.g., 5]
$t_{\rm acc,sh}(E)\approx\frac{3}{u_{1}-u_{2}}\left(\frac{D_{1}(E)}{u_{1}}+\frac{D_{2}(E)}{u_{2}}\right),$ (5.1)
where $u_{1,2}$ and $D_{1,2}$ are the upstream/downstream flow speeds and spatial diffusion coefficients, respectively. For a strong shock with compression ratio $r=u_{1}/u_{2}\simeq 4$ and comparable diffusion coefficients on both sides ($D_{1}\sim D_{2}\sim D$), Eq. (5.1) reduces to the commonly used estimate
$t_{\rm acc,sh}(E)\sim\xi\,\frac{D(E)}{u_{s}^{2}},\qquad\xi\sim\mathcal{O}(10),$ (5.2)
where $\xi$ absorbs order-unity factors depending on the shock structure. In the Bohm-like limit, the diffusion coefficient is
$D_{\rm B}(E)\equiv\frac{1}{3}r_{L}(E)c,\qquad r_{L}(E)=\frac{pc}{eB}\approx\frac{E}{eB},$ (5.3)
for relativistic particles with $E\approx pc$. Combining Eqs. (5.2)–(5.3) yields
$t_{\rm acc,sh}(E)\sim\frac{\xi}{3}\,\frac{r_{L}(E)c}{u_{s}^{2}}\sim\frac{\xi}{3}\,\frac{E}{eB}\,\frac{c}{u_{s}^{2}}.$ (5.4)
To compare the characteristic acceleration timescales of DSA and stochastic acceleration, we form the ratio
$\frac{t_{\rm acc,2}}{t_{\rm acc,sh}}\sim\frac{3}{\xi}\,\left(\frac{c}{v_{\rm tur}}\right)^{2}\left(\frac{u_{s}}{c}\right)^{2}\left(\frac{c}{r_{L}\,\nu_{\rm eff,B}}\right),$ (5.5)
where $\xi\sim\mathcal{O}(10)$ parameterizes order-unity factors in the DSA timescale. The final factor, $c/(r_{L}\nu_{\rm eff,B})$, measures the efficiency of pitch-angle scattering over a gyro-orbit; in the Bohm-like limit, $\nu_{\rm eff,B}\sim c/r_{L}$, this factor is of order unity.
For structure-formation environments, the relevant characteristic velocities are $v_{\rm tur}\sim\epsilon c_{s}$ with $c_{s}\sim 10^{2}-10^{3}~{\rm km\,s^{-1}}$ and $\epsilon\sim 0.1$, implying $v_{\rm tur}/c\sim 10^{-4}-10^{-3}$. Large-scale structure shocks typically have $u_{s}\sim 10^{3}~{\rm km\,s^{-1}}$, corresponding to $u_{s}/c\sim 3\times 10^{-3}$. Adopting Bohm-like scattering ($c/(r_{L}\nu_{\rm eff,B})\sim 1$), these scalings imply
$\frac{t_{\rm acc,2}}{t_{\rm acc,sh}}\sim 10-10^{3},$ (5.6)
up to order-unity factors. Thus, even under optimistic assumptions for the turbulence-driven stochastic mechanism, second-order acceleration is typically one to three orders of magnitude slower than shock acceleration in the same environment.
Consequently, once nonlinear structure formation generates shocks with $u_{s}\sim 10^{3}~{\rm km\,s^{-1}}$ and simultaneously amplifies the magnetic field, DSA operates on substantially shorter timescales. Even if the instability-assisted stochastic channel formally turns on at $z\lesssim z_{\rm on}$, shock acceleration is expected to dominate the production of CRs in that epoch. This supports our central conclusion: a significant pre-shock CR population is unlikely, and efficient CR production is naturally tied to the onset of structure-formation shocks.
6 Summary and Discussion
We have investigated whether stochastic (second-order Fermi) acceleration associated with pressure-anisotropy-driven magnetogenesis can generate a dynamically significant population of cosmic rays (CRs) prior to the onset of nonlinear structure formation. By combining analytic acceleration-time constraints with a Fokker–Planck description of particle transport, we quantified both the conditions under which stochastic acceleration becomes cosmologically viable and the resulting ion momentum distribution in the pre-structure intergalactic medium.
An analytic comparison between the acceleration timescale and the Hubble time yields a critical magnetic field $B_{\rm crit}(z)$ and defines a characteristic turn-on redshift $z_{\rm on}\sim 1.7$, at which stochastic acceleration can in principle begin to compete with cosmological expansion. We find that this turn-on epoch is largely insensitive to uncertainties in the thermal history and microphysical parameters, reflecting the rapid nonlinear amplification of the magnetic field. However, this condition provides only an optimistic upper bound on the efficiency of particle acceleration.
When additional energy loss processes are included, the situation changes qualitatively. In particular, Coulomb interactions in a fully ionized intergalactic medium introduce a strong energy dependence in the effective loss timescale, leading to a refined critical magnetic field $B_{\rm crit}^{\rm eff}(E,z)$. This analysis reveals the existence of an effective injection threshold for stochastic acceleration at energies of order ${\cal O}(10)$ keV, corresponding to $p/(m_{i}c)\sim 5\times 10^{-3}$. Below this scale, Coulomb losses dominate over acceleration ($t_{\rm C}\ll t_{\rm acc}$), efficiently thermalizing particles and suppressing their injection into the acceleration process. At higher energies, where $t_{\rm C}\gg t_{H}$, the acceleration condition approaches the Hubble-limited regime, and the maximum attainable energy remains essentially unchanged from the adiabatic-only estimate.
To quantify the resulting particle population, we solved a Fokker–Planck equation for the isotropic ion distribution over the redshift interval $z=10\rightarrow z_{\rm on}$. The numerical results show that, even under optimistic assumptions corresponding to the strong-scattering limit, the ion distribution remains close to a cooling Maxwellian throughout this epoch. Adiabatic expansion dominates the evolution, and Coulomb losses further suppress the formation of a suprathermal tail by efficiently removing particles near the injection scale. Even when a pre-existing suprathermal component is included in the initial condition, it is rapidly thermalized at low energies and does not lead to a significant enhancement of the nonthermal population.
These results imply that stochastic acceleration during pressure-anisotropy-driven magnetogenesis is unable to generate a dynamically important CR population in the pre-structure universe. In particular, Coulomb losses transform the acceleration problem from a purely cosmological constraint into an energy-dependent injection problem, strongly limiting the fraction of particles that can participate in stochastic energization. As a consequence, the resulting suprathermal population remains energetically subdominant and is unlikely to have a measurable impact on the thermal history of the intergalactic medium or on observable high-energy backgrounds.
Our findings therefore support a picture in which efficient CR production is intrinsically linked to the onset of nonlinear structure formation. While instability-driven scattering can enhance particle transport and may provide a modest level of pre-acceleration, the emergence of a substantial nonthermal population requires the presence of large-scale shocks and sufficiently amplified magnetic fields. The low-level suprathermal population obtained here may serve as a seed for subsequent diffusive shock acceleration (DSA), but its contribution to the overall CR energy budget is expected to be limited. A detailed kinetic study of particle injection and re-acceleration at structure-formation shocks, incorporating the effects of pre-existing turbulence and seed populations, is left for future work.
Acknowledgments
This work was carried out with the support of the Korea Astronomy and Space Science Institute (KASI). The author thanks the referees for their careful reading of the manuscript and for their constructive comments and suggestions, which have significantly improved the quality and clarity of this work.